### v1

# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

Abstract
Sorting is one of the most fundamental algorithms in computer science. Recently, Learned Sorts, which use machine learning to improve sorting speed, have attracted attention. While existing studies show that Learned Sort is experimentally faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity. We propose PCF Learned Sort, a theoretically guaranteed Learned Sort algorithm. We prove that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. We also confirm experimentally that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets. This is the first study to theoretically support the experimental success of Learned Sort, and provides evidence for why Learned Sort is fast.
sorting, learning-augmented algorithms, algorithms with predictions
[…]
do
1 Introduction
The sorting algorithm is one of the most important computer algorithms and has been studied for a long time. Recently, a novel sorting method called Learned Sort has been proposed (Kraska et al., 2019). Learned Sort uses a model that predicts the cumulative density function (CDF) to predict the position of each element in the sorted array, and then uses this prediction to quickly sort the array. Experimental results show that Learned Sort is faster than classical sorting algorithms, including highly tuned counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
On the other hand, there are few theoretical guarantees about the computational complexity of Learned Sort. The first proposed Learned Sort algorithm (Kraska et al., 2019) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity isn’t discussed. The more efficient Learned Sort algorithms proposed later (Kristo et al., 2020, 2021) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications). The goal of this paper is to provide a Learned Sort that is theoretically guaranteed to be computationally efficient.
We propose PCF Learned Sort, which can sort with expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. Furthermore, we prove that PCF Learned Sort can sort with worst-case computational complexity $\mathcal{O}(n\log n)$. Then, we experimentally confirmed that our Learned Sort can sort with the average complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
In Section 2, we first present related work. Then, in Section 3, we introduce our PCF Learned Sort and give the theorems and brief explanations of their proofs that guarantee its computational complexity. Then, in Section 4, we verify our theorems experimentally. We then discuss the limitations and possible future work in Section 5. Finally, in Section 6, we summarize our results and conclusions.
2 Related Work
Our research is in the context of algorithms with predictions, in particular, the sorting algorithms with predictions (Section 2.1). There are two types of sorting algorithms, comparison sorts (Section 2.2) and distribution sorts (Section 2.3), and our proposed method is a distribution sort. However, the idea, implementation, and proof of computational complexity of our method are similar to those of sample sort, which is a type of comparison sort. Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section 2.4).
2.1 Algorithms with predictions
Our research is in the area of algorithms with predictions (Mitzenmacher & Vassilvitskii, 2022). These studies focus on improving the performance (e.g., speed and memory usage) of algorithms for problems such as ski rental (Purohit et al., 2018; Gollapudi & Panigrahi, 2019; Shin et al., 2023), caching (Narayanan et al., 2018; Rohatgi, 2020; Lykouris & Vassilvitskii, 2021; Im et al., 2022), scheduling (Gollapudi & Panigrahi, 2019; Lattanzi et al., 2020; Lassota et al., 2023), and matching (Antoniadis et al., 2023; Dinitz et al., 2021; Sakaue & Oki, 2022) by using predictions from oracles such as machine learning models. The main idea of these novel algorithms is to exploit the characteristics of the data by actively exploiting information not explicitly used by classical algorithms.
In particular, our work is based on research about sorting with predictions. (Kraska et al., 2019) proposed a method for fast sorting that uses a model $\tilde{F}(q)$ trained to approximate the cumulative distribution function (CDF) $F(q)$. This first Learned Sort roughly sorts the array by placing each element $x$ of the array on the $nF(x)$-th entry of the output array. It then sorts the array completely by refining the roughly sorted array using insertion sort. This algorithm has a best case complexity of $\mathcal{O}(n)$, but a worst case complexity of $\mathcal{O}(n^{2})$ because of the exception handling for collisions and the last refining.
Subsequently, implementations have been proposed that improve cache efficiency (Kristo et al., 2020) and are robust against key duplication (Kristo et al., 2021). These also use insertion sorting in the last refinement, resulting in a worst-case complexity of $\mathcal{O}(n^{2})$. The worst-case complexity can be reduced to $\mathcal{O}(n\log n)$ by using a sorting algorithm with small worst-case complexity, such as Introsort (Musser, 1997) or TimSort (McIlroy, 1993). However, this is the same complexity as many effective classical sorting algorithms, and thus does not explain the experimental success of Learned Sort. Some studies have focused on efficiently exploiting the predictions of the oracle, considering the oracle as a black box and not focusing on its inner details (Bai & Coester, 2023). This analysis does not take into account the computational complexity of training and inferring the oracle, so it cannot be directly applied to guarantee the overall computational complexity of sorting.
2.2 Comparison sorts
Sorting algorithms that use comparisons between keys and require no other information about the keys are called comparison sorts. It can be proved that the worst complexity of a comparison sort is at least $\mathcal{O}(n\log n)$. Commonly used comparison sorting algorithms include quick sort, heap sort, merge sort, and insertion sort. The GNU Standard Template Library in C++ uses Introsort (Musser, 1997), an algorithm that combines quick sort, heap sort, and insertion sort. Java (Java, 2023) and Python up to version 3.10 (Peters, 2002) use TimSort (McIlroy, 1993), an improved version of merge sort. Python 3.11 and later use Powersort (Munro & Wild, 2018), a merge sort that determines a near-optimal merge order.
Sample sort (Frazer & McKellar, 1970) is an extension of quick sort that uses multiple pivots, whereas quick sort uses only one pivot. Sample sort samples a small number of keys from the array, determines multiple pivots, and uses them to partition the array into multiple buckets. The partitioning is repeated recursively until the array is sufficiently small. The in-place parallel superscalar sample sort (Axtmann et al., 2022) is one of the most efficient sample sort implementation. They theoretically guarantee its computational and cache efficiency by proving a theorem about the probability that the pivots partition the array (nearly) equally.
2.3 Distribution sorts
Distribution sorts use information other than key comparison. Radix sort and counting sort are the most common types of distribution sorts. Radix sort uses counting sort as a subroutine for each digit. When the number of digits in the array element is $w$, the computational complexity of radix sort is $\mathcal{O}(wn)$. Thus, radix sort is particularly effective when the number of digits is small. There are several variants of radix sort, such as SkaSort (Skarupke, 2016), which is optimized using a technique called “American flag sort,” and RegionSort (Obeya et al., 2019), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
2.4 Learned Index
Kraska et al. showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models, and named such novel data structures Learned Index (Kraska et al., 2018). Since then, various learning augmented B-trees (Wang et al., 2020; Kipf et al., 2020) and learning augmented Bloom filters (Mitzenmacher, 2018; Vaidya et al., 2021; Sato & Matsui, 2023) have been proposed. There are several researches on learning augmented B-trees whose performance is theoretically guaranteed. PGM-index (Ferragina & Vinciguerra, 2020) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$. (Zeighami & Shahabi, 2023) proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution.
3 Methods
Figure 1: PCF Learned Sort: First, the input array is partitioned into $\gamma+1$ buckets using a CDF model-based method. Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a “standard” sorting algorithm of complexity $\mathcal{O}(n\log n)$ (e.g. IntroSort). Otherwise, the recursive model-based bucketing is repeated. Finally, the sorted arrays are concatenated. The CDF model used for bucketing is a Piecewise Constant Function (PCF). The function is constant within each interval and the interval widths are constant.
Algorithm 1 The Learned Sort algorithm
[…]
This section describes our PCF Learned Sort. First, in Section 3.1, we give an overview of the algorithm and the key theorems used to guarantee the expected and worst-case computational complexity. Next, in Section 3.2, we introduce the specific algorithm for PCF Learned Sort and give a guarantee on its computational complexity.
3.1 Method Overview
Our algorithm basically repeats recursive model-based bucketing until the array is small enough or the bucketing “fails.” In model-based bucketing, a CDF model is trained and then used to partition an array of length $n$ into approximately $\gamma$ buckets, satisfying the following two conditions: the elements in $i$-th bucket are smaller than any elements in $(i+1)$-th bucket, and the buckets have approximately the same size with high probability. If we set $\gamma=n^{c}$ (where $10<c<$ 1 is a constant) and the partitioning into buckets is done well, the size of each bucket will be approximately $n^{1-c}$, so the size of the bucket in the $k$-th recurrence depth will be approximately $n^{(1-c)^{k}}$. Here,
$k\geq-\frac{\log\log n}{\log(1-c)}\Rightarrow n^{(1-c)^{k}}$ $\leq n^{(1-c)^{-\frac{\log\log n}{\log(1-c)}}}$ (1)
$=n^{\frac{1}{\log n}}$ (2)
$=e,$ (3)
where $\log$ is the natural logarithm, so the base is $e$. That is, with a recursion depth of $k=\Omega(\log\log n)$, the size of bucket is small enough, i.e., the max recursion depth is $\mathcal{O}(\log\log n)$. Thus, if the expected computational complexity of the partition is $\mathcal{O}(n)$ and the probability that the bucketing “fails” is sufficiently small, the total computational complexity is $\mathcal{O}(n\log\log n)$. Under mild assumptions on the distribution, we can prove that this fast and accurate partition can be achieved by using the Piecewise Constant Function (PCF) as the CDF model.
Let $\mathcal{D}~{}(\subset{\mathbb{R}})$ be the range of possible values of the input array and $\bm{x}~{}(\in\mathcal{D}^{n})$ be the input array. If the length of the input array, denoted as $|\bm{x}|$, is less than $\tau$, our algorithm sorts the input array using a “standard” sorting method with $\mathcal{O}(n\log n)$ complexity, such as IntroSort, where $\tau$ is a predetermined constant. Otherwise, if $|\bm{x}|\geq\tau$, model-based bucketing is performed.
The model-based bucketing method $\mathcal{M}$ takes an input array $\bm{x}$ and partitions it into several buckets. First, it determines the parameter $\gamma~{}(\in\mathbb{N})$, which determines the number of buckets. $\gamma$ is determined as a function of the array length $n$. The number of buckets returned by $\mathcal{M}$ is $\gamma+1$. Next, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\mapsto[0,1]$. The $\tilde{F}(q)$ is trained to predict the empirical cumulative distribution function of $\bm{x}$ defined as:
$F_{\bm{x}}(q)\coloneqq\frac{|\{i\in\{1,\dots,n\}\mid x_{i}\leq q\}|}{|\bm{x}|}.$ (4)
The specific model and training method is explained in Section 3.2. Finally, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets. All $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$, are initialized to be empty, and then for each $i\in\{1,\dots,n\}$, $x_{i}$ is appended to $\bm{c}_{\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1}$. This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., |$|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
The CDF model $\tilde{F}$ must be a non-decreasing function to guarantee the computational complexity of our algorithm. In this case, the bucket with the larger ID gets the larger value, i.e.,
$p\in\bm{c}_{j},q\in\bm{c}_{k},1\leq j<k\leq\gamma+1\Rightarrow p<q.$ (5)
This means that each bucket is responsible for a disjoint and continuous interval. Let $t_{j}=\min_{x\in\bm{c}_{j}}x~{}(j=1,\dots,\gamma+1),~{}t_{\gamma+2}=\infty$, then the $j$-th bucket $\bm{c}_{j}$ ($j=1,\dots,\gamma+1$) is responsible for a continuous interval $\mathcal{I}_{j}\coloneqq[t_{j},t_{j+1})$.
After model-based bucketing, our algorithm determines for each bucket whether the bucketing was “successful” or not. For each $j=1,\dots,\gamma+1$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$, where $\delta$ is an integer determined by $n$. If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the “standard” sorting method (e.g., IntroSort). If $|\bm{c}_{j}|<\delta$ (which means the bucketing “succeeds”), the bucket is sorted by recursively calling our Learned Sort algorithm. Note that the parameters such as $\gamma$ and $\delta$ are redetermined according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket. After each bucket is sorted, the buckets are concatenated. The input array $\bm{x}$ is sorted by the above procedure. The basic idea of our algorithm is visualized in Figure 1, and the pseudocode is given in Algorithm 1.
The following is a lemma about the worst-case complexity of our Learnd Sort as defined above.
Lemma 3.1.
Assume that there exists an model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$. Also, assume that the “standard” sort algorithm used for the too large or too small buckets has a worst-case complexity of $\mathcal{O}(n\log n)$. The worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $10<d<$ 1) is $\mathcal{O}(n\log n)$.
This lemma can be intuitively shown from the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) each element of the input array $\bm{x}$ undergoes several bucketing and only one “standard” sort. (i) can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$, and (ii) is evident from the algorithm’s design since the buckets sorted by “standard” sort are now left only to be concatenated. The exact proof is given in Section A.1.
Next, we introduce an important lemma about the expected complexity of our Learned Sort. The following assumption is necessary to guarantee the expected complexity.
Assumption 3.2.
The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\mapsto\mathbb{R}_{\geq 0}$.
We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\mapsto\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subset\mathcal{D}$, i.e.,
[…]
The expected computational complexity of our proposed Learned Sort is guaranteed by the following lemma.
Lemma 3.3.
Let $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ be the array formed by sampling $n$ times independently according to $f_{\mathcal{I}}(x)$. Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~{}(\in(0,1))$ that satisfy the following for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
  • $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$, with an expected complexity of $\mathcal{O}(n)$. That is, $\mathcal{M}$ can take the array $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ as input and divide it into $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{j=\gamma+1}$ satisfying Equation 5, with an expected complexity of $\mathcal{O}(n)$.
  • $\gamma+1=\mathcal{O}(n)$.
  • $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$.
The expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
This lemma can be proved intuitively by the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$. (i) is the same as in the explanation of the proof of Lemma 3.1. (ii) can be shown from the fact that the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases. See Section A.2 for the exact proof.
Note that the assumption of Lemma 3.3 includes “$\mathcal{M}$ works well with high probability for any $\mathcal{I}$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing. The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~{}(\subset\mathcal{D})$.
3.2 PCF Learned Sort
We propose PCF Learned Sort as an implementation that satisfies the assumptions of Lemma 3.1 and Lemma 3.3, and thus has a guarantee that the worst-case complexity is $\mathcal{O}(n\log n)$ and the expected complexity is $\mathcal{O}(n\log\log n)$. PCF Learned Sort approximates the CDF by a Piecewise Constant Function (PCF). Piecewise constant function is a function that has intervals of equal width and outputs the same value in each interval (the right side of Figure 1). The study that develops a Learned Index with a theoretical guarantee on its complexity (Zeighami & Shahabi, 2023) also used PCF as a CDF model.
The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ in the following way. First, the parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, the length of the input array. The parameter $\alpha$ is the number of samples to train the CDF model, and $\beta$ is the number of intervals in the PCF. Next, the piecewise constant function is trained by counting the number of samples in each interval. Using $x_{\mathrm{min}}=\min_{i}{x_{i}}$ and $x_{\mathrm{max}}=\max_{i}{x_{i}}$, let $i(x)$ be a function defined as follows:
$i(x)=\left\lfloor\frac{x-x_{\mathrm{min}}}{x_{\mathrm{max}}-x_{\mathrm{min}}}\beta\right\rfloor+1.$ (7)
$\alpha$ samples are taken at random from $\bm{x}$ to form $\bm{a}~{}(\in\mathcal{D}^{\alpha})$, and then $i(x)$ is used to form the array $\bm{b}~{}(\in{\mathbb{Z}_{\geq 0}}^{\beta+1})$ defined as follows:
$b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|.$ (8)
This counting is corresponding to the training of the PCF. Note that $\bm{b}$ is an non-decreasing non-negative array and $b_{\beta+1}=\alpha$, i.e., $0\leq b_{1}\leq b_{2}\leq\dots\leq b_{\beta+1}=\alpha$.
The inference for the CDF model $\tilde{F}(x)$ is performed using $i(x)$, $\bm{b}$, and the following equation:
$\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}.$ (9)
Since $i(x)$ is a non-decreasing function and $\bm{b}$ is also a non-decreasing array, $\tilde{F}(x)$ is a non-decreasing function. Also, $0\leq\tilde{F}(x)\leq 1$ because $0\leq b_{i}\leq\alpha$ for every $i$.
The following is a lemma to bound the probability that $\mathcal{M}_{\mathrm{PCF}}$ will “fail” bucketing. This lemma is important to guarantee the expected computational complexity of PCF Learned Sort.
Lemma 3.4.
Let $\sigma_{1}$ and $\sigma_{2}$ be respectively the lower and upper bounds of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$. That is, $x\in\mathcal{D}\Rightarrow\sigma_{1}\leq f(x)\leq\sigma_{2}$.
Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
$\begin{split}&K\geq 1\\ \Rightarrow&\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},\end{split}$ (10)
where
$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}.$ (11)
The proof of this lemma is based on and combines proofs from two existing studies. The first is Lemma 5.2. from a study of IPS4(Axtmann et al., 2022), an efficient sample sort. This lemma guarantees the probability of a “successful recursion step” when selecting pivots from samples and using them to perform a partition. Although the definition of “successful recursion step” in this paper is similar to our definition of “successful bucketing” (more precisely, in the special case of $\delta=3n/\gamma$), this lemma is for the method that do not use the CDF model, so the proof cannot be applied directly to our case. Another proof we refer to is the proof of Lemma 4.5. from a study that addressed the computational complexity guarantee of the Learned Index (Zeighami & Shahabi, 2023). This lemma provides a probabilistic guarantee for the error between the output of the PCF and the empirical CDF. Some modifications are required to adapt it to the context of sorting and to attribute it to the probability of bucketing failure, i.e., $\Pr[\exists j,|\bm{c}_{j}|>\delta]$. By appropriately combining the proofs of these two lemmas, Lemma 3.4 can be proved. The exact proof is given in Section A.3.
Using Lemma 3.1, Lemma 3.3 and Lemma 3.4, we can prove the following theorems.
Theorem 3.5.
If we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
Proof.
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$. This is because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element. Therefore, from Lemma 3.1, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$. ∎
Theorem 3.6.
Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
Then, if we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
Proof.
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$.
Since $K=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K\geq 1$ for sufficiently large $n$, and
$\begin{split}\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}&=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\\ &\leq\mathcal{O}\left(\frac{1}{\log n}\right).\end{split}$ (12)
Therefore, from Lemma 3.3, the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort, since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$ and $\delta$, are determined without the knowledge. In other words, PCF Learned Sort can sort in expected $\mathcal{O}(n\log\log n)$ complexity as long as $0<\sigma_{1}\leq f(x)\leq\sigma_{2}<\infty$, even if it does not know the exact value of $\sigma_{1}$ and $\sigma_{2}$. If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem 3.6, i.e., $\sigma_{1}=0$ or $\sigma_{2}=\infty$, then the expected complexity of PCF Learned Sort is increased to $\mathcal{O}(n\log n)$, but from Theorem 3.5, it cannot be greater than $\mathcal{O}(n\log n)$.
4 Experiments
In this section, we confirm our theorems experimentally. First, we confirm that the computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real data. Then, we conduct experiments with various parameter settings and experimentally confirm Lemma 3.4, a lemma that bounds the probability of bucketing failure and is an important lemma to guarantee the expected computational complexity of PCF Learned Sort.
4.1 Results on Synthetic Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Number of operations to sort the synthetic array. Below each graph is a histogram that visualizes the distribution of each dataset.
We experimented with datasets created from the following four types of synthetic distributions: uniform distribution ($\mathsf{min}=0,\mathsf{max}=1$), normal distribution ($\mu=0,\sigma=1$), exponential distribution ($\lambda=1$), lognormal distribution ($\mu=0,\sigma=1$). We took $n=10^{3},\dots,10^{8}$ samples from each distribution independently, and meticulously counted the total number of basic operations to sort the sampled array. Here, the basic operations consists of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access. We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware (this is the same idea as the metric selection in the experiment of  (Zeighami & Shahabi, 2023)). The parameters of PCF Learned Sort are set as in Theorem 3.5 and Theorem 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$. As the “standard” sort algorithm used in PCF Learned Sort, we used quick sort, which is known to have an expected computational complexity of $\mathcal{O}(n\log n)$. We then compared the number of operations between PCF Learned Sort and the (plain) quick sort.
Figure 2 shows the number of operations divided by the length of input array, $n$. It shows the mean and standard deviation of the $10$ measurements for each condition. We see that our PCF Learned Sort has up to 2.8 times fewer operations than quick sort. Also, while the graph of quick sort is almost linear, the graph of PCF Learned Sort is almost flat. This suggests that PCF Learned Sort has a computational complexity much smaller than $\mathcal{O}(n\log n)$.
4.2 Results on Real Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Number of operations to sort the real data array. Below each graph is a histogram that visualizes the distribution of each dataset.
We used four real datasets, NYC, Wiki, OSM, and Books, which are described below. NYC: pick-up datetimes in the yellow taxi trip records (Kristo, 2021). Wiki: Wikipedia article edit timestamps (Marcus et al., 2020). OSM: uniformly sampled OpenStreetMap locations represented as Google S2 CellIds (Marcus et al., 2020). Books: book sale popularity data from Amazon (Marcus et al., 2020). For each dataset, we randomly sample $n=10^{3},\dots,10^{8}$ elements, shuffle the sampled elements, and then use it as an input array to examine the performance of the algorithms. As in the case of the synthetic dataset, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and quick sort is used as “standard” sort.
Figure 3 shows plots of the number of operations divided by the length of input array, $n$. Again, the mean and standard deviation of $10$ measurements are shown. We see that our PCF Learned Sort has up to 2.7 times fewer operations than quick sort (as in the case of the synthetic dataset), and that PCF Learned Sort has much less computational complexity than $\mathcal{O}(n\log n)$.
We can also see that the number of operations required for the OSM dataset is about $1.25$ times larger than that for the other datasets. Although not mentioned explicitly, the same tendency is observed in the experiments of existing Learned Sort studies (Kristo et al., 2020, 2021). Our theoretical analysis can explain this phenomenon from the perspective of $\sigma_{2}/\sigma_{1}$, in other words, the strength of the distributional bias. From Equation 11 and Equation 10, the larger $\sigma_{2}/\sigma_{1}$ is, the smaller $K$ is, and the larger (the upper bound of) the probability of bucketing failure. Therefore, the larger $\sigma_{2}/\sigma_{1}$ is, the more the computational complexity will move away from $\mathcal{O}(n\log\log n)$ and closer to the worst-case computational complexity, $\mathcal{O}(n\log n)$. In fact, the histograms of each dataset show that the OSM dataset has sharp peaks and many intervals with almost no data, indicating that it has a distribution with large $\sigma_{2}/\sigma_{1}$.
4.3 Confirmation of Lemma 3.4
Figure 4: Heatmap showing the experimental frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$. The variables $a,b,c,d$, except those on the x- and y-axes, were set to $0.75$. The white dotted line represents the parameters that make the right side of Equation 10 equal to 0.5.
Lemma 3.4 bounds the probability that a bucket of size greater than $\delta$ exists. This is an important lemma that allows us to guarantee the expected computational complexity of PCF Learned Sort. Here, we experimentally confirm that this upper bound is appropriate.
We have experimented with $\alpha=\lfloor n^{a}\rfloor,\beta=\lfloor n^{b}\rfloor,\gamma=\lfloor n^{c}\rfloor,\delta=\lfloor n^{d}\rfloor$, varying $a,b,c,d$ from 0.05 to 0.95 at 0.05 intervals. For each $a,b,c,d$ setting, the following was repeated $100$ times: we took $n=10^{6}$ elements from the uniform distribution to form the input array and divided the array into $\gamma+1$ buckets by $\mathcal{M}_{\mathrm{PCF}}$, and checked whether or not $\exists j,|\bm{c}_{j}|>\delta$. Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the experimental frequency at which bucketing “fails.”
Heat maps in Figure 4 show the experimental frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed and the other two parameters are varied. The value of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure 4 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$. The white dotted line represents the parameter so that the right side of Equation 10 is 0.5. That is, Lemma 3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
We can see that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often. In other words, we can see that the theoretical upper bound from Lemma 3.4 agrees well (to some extent) with the actual probability. We can also confirm that, as Lemma 3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
5 Limitation and Future Work
We proposed PCF Learned Sort, whose expected computational complexity can be proved to be $\mathcal{O}(n\log\log n)$ under the assumption that there exist $0<\sigma_{1}$ and $\sigma_{2}<\infty$ satisfying $\sigma_{1}<f(x)<\sigma_{2}$. Thus, our proof is not applicable to the case of a distribution where there exists $x\in\mathcal{D}$ such that $f(x)=0$ or $f(x)=\infty$. The development of a theory and/or algorithm that is applicable for a wider class of distributions is a future work.
Other future work may include incorporating a more advanced CDF approximation method with a theoretical guarantee into our Learned Sort algorithm. Using a different CDF model and a bucketing algorithm that satisfies the assumptions in Lemma 3.3, we may be able to develop an even more efficient sorting algorithm. In other words, a more accurate and faster CDF approximation method can lead to an improvement of the Learned Sort.
6 Conclusion
We proposed PCF Learned Sort and proved theoretically that its worst-case computational complexity is $\mathcal{O}(n\log n)$ (without assumptions) and its expected computational complexity is $\mathcal{O}(n\log\log n)$ under mild assumptions on the distribution. We then confirm this computational complexity experimentally on both synthetic and real data. Furthermore, we provide a theoretical and intuitive explanation for the slow speed of Learned Sort on some datasets. This is the first study to theoretically support the experimental success of Learned Sort and provides insight into why and when Learned Sort is fast.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs
In this Appendix, we give the proofs omitted in the main paper. In Section A.1, we give the proof of Lemma 3.1, which is important for proving the worst-case complexity of PCF Learned Sort. Section A.2 and Section A.3 give proofs of Lemma 3.3 and Lemma 3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort.
A.1 Proof of Lemma 3.1
Proof.
Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.1 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the worst-case complexity of the “standard” sort algorithm and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, and $\gamma+1=\mathcal{O}(n)$,
$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n.$ (13)
$\exists~{}n_{2},l_{2}~{}(>0),$ $~{}~{}n\geq n_{2}\Rightarrow R(n)\leq l_{2}n.$ (14)
$\exists~{}n_{3},l_{3}~{}(>0),$ $~{}~{}n\geq n_{3}\Rightarrow\gamma+1\leq l_{3}n.$ (15)
First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $P(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations. This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the “standard” sort algorithm.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Let $\mathcal{S}_{\gamma,k}$ be the set consisting of all $(\gamma+1)$-dimensional vectors of positive integers whose sum is $k$, i.e., $\mathcal{S}_{\gamma,k}\coloneq\left\{\bm{s}\in\mathbb{Z}_{\geq 0}^{\gamma}\mid\sum_{i=1}^{\gamma+1}s_{i}=k\right\}$. Then,
$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+ls_{i}\log s_{i}\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\left(C+ls_{i}\log k^{d}\right)\right\}\\ &\leq l_{2}k+C_{1}(\gamma+1)+l_{1}k\log k+C(\gamma+1)+lk\log k^{d}\\ &\leq l_{2}k+C_{1}l_{3}k+l_{1}k\log k+Cl_{3}k+lkd\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}-l(1-d)\right\}k\log k+(C+lk\log k).\end{split}$ (16)
Therefore, if we take $l$ such that
$\frac{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}}{1-d}\leq l,$ (17)
then $P(k)\leq C+lk\log k$ (note that the left side of the following equation is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n\in\mathbb{N}$. ∎
A.2 Proof of Lemma 3.3
Proof.
The proof approach is the same as in Lemma 3.1, but in Lemma 3.3 the “expected” computational complexity is bounded. The following two randomnesses are considered to compute the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.3 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the expected complexity of the “standard” sort algorithm and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n,$ (18)
$\exists~{}n_{2},l_{2}~{}(>0),$ $~{}~{}n\geq n_{2}\Rightarrow R(n)\leq l_{2}n,$ (19)
$\exists~{}n_{3},l_{3}~{}(>0),$ $~{}~{}n\geq n_{3}\Rightarrow\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]\leq\frac{l_{3}}{\log n},$ (20)
$\exists~{}n_{4},l_{4}~{}(>0),$ $~{}~{}n\geq n_{4}\Rightarrow\gamma+1\leq l_{4}n.$ (21)
In the following, we prove $T(n)=\mathcal{O}(n\log\log n)$ by mathematical induction.
First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $T(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
$\begin{split}T(k)&\leq R(k)+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot S(|\bm{c}_{j}|)+\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr\left[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor\right]\cdot\mathbb{E}\left[\sum_{j=1}^{\gamma+1}S(|\bm{c}_{j}|)\right]+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot\{C_{1}(\gamma+1)+l_{1}k\log k\}+\mathbb{E}\left[\sum_{i=1}^{\gamma+1}T(\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|))\right].\end{split}$ (22)
Here, from $k\geq n_{3}$ and $k\geq n_{4}$,
$\begin{split}\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot\{C_{1}(\gamma+1)+l_{1}k\log k\}&\leq\frac{l_{3}}{\log k}\cdot(C_{1}l_{4}k+l_{1}k\log k)\\ &\leq(C_{1}l_{3}l_{4}+l_{1}l_{3})k.\end{split}$ (23)
Also, from the assumption of induction and $k\geq n_{4}$,
$\begin{split}\mathbb{E}\left[\sum_{i=1}^{\gamma+1}T(\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|))\right]&\leq\mathbb{E}\left[\sum_{i=1}^{\gamma+1}\left\{C+l\cdot\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|)\log\log\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|)\right\}\right]\\ &\leq\mathbb{E}\left[C(\gamma+1)+\sum_{i=1}^{\gamma+1}l\cdot|\bm{c}_{j}|\log\log\lfloor k^{d}\rfloor\right]\\ &\leq Cl_{4}k+lk\log\log\lfloor k^{d}\rfloor\\ &\leq Cl_{4}k+lk\log d+lk\log\log k.\end{split}$ (24)
Therefore,
$\begin{split}T(k)&\leq l_{2}k+(C_{1}l_{3}l_{4}+l_{1}l_{3})k+Cl_{4}k+lk\log d+lk\log\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}l_{4}+l_{1}l_{3}+Cl_{4}-l\log\frac{1}{d}\right\}k+(C+lk\log\log k).\end{split}$ (25)
Therefore, if we take $l$ such that
$\frac{l_{2}+C_{1}l_{3}l_{4}+l_{1}l_{3}+Cl_{4}}{\log\frac{1}{d}}\leq l,$ (26)
then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation 26 is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
A.3 Proof of Lemma 3.4
To prove Lemma 3.4, we first present the following lemma.
Lemma A.1.
Let $\bm{e}~{}(\in\mathcal{I}^{n})$ be a sorted version of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ and $\Delta\coloneqq(x_{\mathrm{max}}-x_{\mathrm{min}})/\beta$ (where $x_{\mathrm{min}}$ and $x_{\mathrm{max}}$ are the minimum and maximum values of $\bm{x}_{\mathcal{I}}$, respectively).
Also, define the set $\mathcal{S}_{r},\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (27)
Using this definition, define $Y_{jr},Z_{jr},Y_{r},Z_{r}$ as follows ($j=1,\dots,\alpha,~{}r=1,\dots,n$):
$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~{}~{}~{}~{}~{}~{}~{}~{}Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},$ (28)
$Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~{}~{}~{}~{}~{}~{}~{}~{}Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}.$ (29)
If the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, then the following holds:
$\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\lor~{}\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right).$ (30)
Proof.
We prove the contraposition of the lemma. That is, we prove that $e_{r}$ is allocated to a bucket smaller than $\delta$ under the assumption that $\left(r<\frac{\delta}{2}+1~{}\lor~{}Y_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\land~{}\left(r>n-\frac{\delta}{2}~{}\lor~{}Z_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)$.
For convenience, we hypothetically define $e_{0}=-\infty,e_{n+1}=\infty$, and assign $e_{0}$ to the 0th bucket and $e_{n+1}$ to the $(\gamma+2)$-th bucket. The size of the 0th bucket and the $(\gamma+2)$-th bucket are both 1.
First, we prove that $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}$ are assigned to different buckets. When $n-\delta/2<r\leq n$, $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}=e_{n+1}$ are obviously assigned to different buckets. When $r\leq n-\delta/2$, the ID of the bucket to which $e_{r}$ is assigned is
$\begin{split}\lfloor\tilde{F}(e_{r})\gamma\rfloor+1&=\left\lfloor\frac{\gamma}{\alpha}b_{i(e_{r})}\right\rfloor+1\\ &\leq\frac{\gamma}{\alpha}b_{i(e_{r})}+1\\ &=\frac{\gamma}{\alpha}\left|\{j\mid i(a_{j})\leq i(e_{r})\}\right|+1\\ &\leq\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|+1.\end{split}$ (31)
The ID of the bucket to which $e_{\min\left(r+\delta/2,n+1\right)}=e_{r+\delta/2}$ is assigned is
$\begin{split}\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1&=\left\lfloor\frac{\gamma}{\alpha}b_{i(e_{r+\delta/2})}\right\rfloor+1\\ &>\frac{\gamma}{\alpha}b_{i(e_{r+\delta/2})}\\ &=\frac{\gamma}{\alpha}\left|\{j\mid i(a_{j})\leq i(e_{r+\delta/2})\}\right|\\ &\geq\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|.\end{split}$ (32)
Thus, taking the difference between these two bucket IDs,
$\begin{split}\left(\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1\right)-\left(\lfloor\tilde{F}(e_{r})\gamma\rfloor+1\right)&>\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|-\left(\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|+1\right)\\ &=\frac{\gamma}{\alpha}\left|\{j\mid e_{r}+\Delta<a_{j}\leq e_{r+\delta/2}-\Delta\}\right|-1\\ &=\frac{\gamma}{\alpha}\left|\mathcal{T}_{r}\right|-1\\ &=\frac{\gamma}{\alpha}\sum_{j=1}^{\alpha}Z_{jr}-1\\ &=\frac{\gamma}{\alpha}Z_{r}-1\\ &\geq 0.\end{split}$ (33)
Therefore, $e_{r}$ and $e_{\min\left(r+\delta/2,n+1\right)}$ are assigned to different buckets.
In the same way, we can prove that $e_{\max(0,r-\delta/2)}$ and $e_{r}$ are also assigned to different buckets. Thus, the size of the bucket to which $e_{r}$ is assigned is at most $\delta-1$ (at most from $e_{\max(0,r-\delta/2)+1}$ to $e_{\min\left(r+\delta/2,n+1\right)-1}$), and the contraposition of the lemma is proved. ∎
Using Lemma A.1, we can prove Lemma 3.4.
Proof.
[…]
Thus, when $r\geq\frac{\delta}{2}+1$,
$\begin{split}\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}_{r}\right|\right]&=\mathbb{E}\left[\frac{\delta}{2}-\left|\{k\mid e_{r-\delta/2}+\Delta<e_{k}\leq e_{r}-\Delta\}\right|\right]\\ &=\mathbb{E}\left[\left|\{k\mid e_{r-\delta/2}<e_{k}\leq e_{r-\delta/2}+\Delta\}\right|\right]+\mathbb{E}\left[\left|\{k\mid e_{r}-\Delta<e_{k}\leq e_{r}\}\right|\right]\\ &\leq nq+nq\\ &\leq\frac{2\sigma_{2}n}{\sigma_{1}\beta}.\end{split}$ (35)
Thus, when $r\geq\frac{\delta}{2}+1$,
$\begin{split}\mathbb{E}[Y_{r}]&=\frac{\alpha}{n}\mathbb{E}\left[\left|\mathcal{S}_{r}\right|\right]\\ &=\frac{\alpha}{n}\left(\frac{\delta}{2}-\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}_{r}\right|\right]\right)\\ &\geq\frac{\alpha\delta}{2n}-\frac{2\sigma_{2}\alpha}{\sigma_{1}\beta}\\ &=\frac{\alpha K}{\gamma}.\end{split}$ (36)
Here, when $K\geq 1$, we have
$10\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<$ 1 . (37)
Therefore, from the Chernoff bound,
$\begin{split}\Pr\left[Y_{r}\leq\frac{\alpha}{\gamma}\right]&=\Pr\left[Y_{r}\leq\left\{1-\left(1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}\right)\right\}\mathbb{E}\left[Y_{r}\right]\right]\\ &\leq\exp\left\{-\frac{1}{2}\left(1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}\right)^{2}\mathbb{E}\left[Y_{r}\right]\right\}\\ &\leq\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (38)
In the same way, we can prove that when $r\leq n-\frac{\delta}{2}$,
$\Pr\left[Z_{r}\leq\frac{\alpha}{\gamma}\right]\leq\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (39)
Thus, by defining $E_{r}$ to be the event “$e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma A.1,
$\begin{split}\Pr[E_{r}]&\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~{}\lor~{}\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~{}\land~{}Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]+\Pr\left[\left(r\leq n-\frac{\delta}{2}~{}\land~{}Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\begin{cases}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&\left(r<\frac{\delta}{2}+1~{}\lor~{}r>n-\frac{\delta}{2}\right)\\ 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&(\mathrm{else})\end{cases}\\ &\leq 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (40)
Therefore,
$\mathbb{E}\left[\sum_{r=1}^{n}\mathbbm{1}[E_{r}]\right]\leq 2n\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}$ (41)
Then, noting that the number of buckets with size greater than or equal to $\delta$ is less than or equal to $\sum_{r=1}^{n}\mathbbm{1}[E_{r}]/\delta$,
$\mathbb{E}\left[|\{j\mid|\bm{c}_{j}|>\delta\}|\right]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (42)
Then, from Markov’s inequality, we have
$\begin{split}\Pr[\exists j,|\bm{c}_{j}|>\delta]&=\Pr[|\{j\mid|\bm{c}_{j}|>\delta\}|\geq 1]\\ &\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (43)

### v3

@@
 Recently, Learned Sorts, which use machine learning to improve sorting speed, have attracted attention.
-While existing studies show that Learned Sort is experimentally faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity.
-We propose PCF Learned Sort, a theoretically guaranteed Learned Sort algorithm.
+While existing studies show that Learned Sort is empirically faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity.
+We propose Piecewise Constant Function (PCF) Learned Sort, a theoretically guaranteed Learned Sort algorithm.
 We prove that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution.
-We also confirm experimentally that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
-This is the first study to theoretically support the experimental success of Learned Sort, and provides evidence for why Learned Sort is fast.
-sorting, learning-augmented algorithms, algorithms with predictions
-[…]
-do
+We also confirm empirically that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
+This is the first study to theoretically support the empirical success of Learned Sort, and provides evidence for why Learned Sort is fast.
 1 Introduction
-The sorting algorithm is one of the most important computer algorithms and has been studied for a long time.
-Recently, a novel sorting method called Learned Sort has been proposed (Kraska et al., 2019).
-Learned Sort uses a model that predicts the cumulative density function (CDF) to predict the position of each element in the sorted array, and then uses this prediction to quickly sort the array.
-Experimental results show that Learned Sort is faster than classical sorting algorithms, including highly tuned counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
-On the other hand, there are few theoretical guarantees about the computational complexity of Learned Sort.
-The first proposed Learned Sort algorithm (Kraska et al., 2019) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity isn’t discussed.
-The more efficient Learned Sort algorithms proposed later (Kristo et al., 2020, 2021) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications).
-The goal of this paper is to provide a Learned Sort that is theoretically guaranteed to be computationally efficient.
-We propose PCF Learned Sort, which can sort with expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution.
-Furthermore, we prove that PCF Learned Sort can sort with worst-case computational complexity $\mathcal{O}(n\log n)$.
-Then, we experimentally confirmed that our Learned Sort can sort with the average complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
-In Section 2, we first present related work.
-Then, in Section 3, we introduce our PCF Learned Sort and give the theorems and brief explanations of their proofs that guarantee its computational complexity.
-Then, in Section 4, we verify our theorems experimentally.
-We then discuss the limitations and possible future work in Section 5.
-Finally, in Section 6, we summarize our results and conclusions.
+Sorting is one of the most fundamental algorithms in computer science and has been extensively studied for many years.
+Recently, a novel sorting method called Learned Sort has been proposed (kraska2021sagedb).
+In Learned Sort, a machine learning model is trained to estimate the distribution of elements in the input array, specifically, the cumulative distribution function (CDF).
+The model’s predictions are then used to rearrange the elements, followed by a minor refinement step to complete the sorting process.
+Empirical results show that Learned Sort is faster than classical sorting algorithms, including highly optimized counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
+On the other hand, there are few theoretical guarantees regarding the computational complexity of Learned Sort.
+The first proposed Learned Sort algorithm (kraska2021sagedb) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity is not discussed.
+The more efficient Learned Sort algorithms proposed later (kristo2020case; kristo2021defeating) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications).
+The goal of this paper is to develop a Learned Sort that is theoretically guaranteed to be computationally efficient.
+We propose Piecewise Constant Function (PCF) Learned Sort, which can sort with an expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution.
+In addition, we show that PCF Learned Sort admits a worst-case complexity guarantee that depends on the choice of the internal sorting algorithm.
+For instance, if MergeSort—whose worst-case complexity is $\mathcal{O}(n\log n)$—is used, then the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
+If we instead use the algorithm by han2020sorting, which has a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$, then the worst-case complexity of PCF Learned Sort becomes $\mathcal{O}(n\sqrt{\log n})$.
+We then empirically confirm that our Learned Sort can sort with a complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
+Independently and concurrently, zeighami2024theoretical explored complexity-guaranteed Learned Sort.
+While our PCF Learned Sort is motivated by similar principles and incorporates comparable design choices, several key distinctions exist between our approach and theirs.
+A comprehensive comparison between our method and that of zeighami2024theoretical is provided in Section˜5.
+This paper is organized as follows.
+Section˜2 reviews related literature.
+Section˜3 introduces PCF Learned Sort and provides complexity theorems with proof sketches.
+Section˜4 provides empirical validation of the theoretical results.
+We discuss implications and limitations in Section˜5, and conclude in Section˜6.
 2 Related Work
-Our research is in the context of algorithms with predictions, in particular, the sorting algorithms with predictions (Section 2.1).
-There are two types of sorting algorithms, comparison sorts (Section 2.2) and distribution sorts (Section 2.3), and our proposed method is a distribution sort.
+Our research is in the context of algorithms with machine learning (Section˜2.1).
+There are two types of sorting algorithms, comparison sorts (Section˜2.2) and non-comparison sorts (Section˜2.3), and our proposed method is a non-comparison sort.
 However, the idea, implementation, and proof of computational complexity of our method are similar to those of sample sort, which is a type of comparison sort.
-Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section 2.4).
-2.1 Algorithms with predictions
-Our research is in the area of algorithms with predictions (Mitzenmacher & Vassilvitskii, 2022).
-These studies focus on improving the performance (e.g., speed and memory usage) of algorithms for problems such as ski rental (Purohit et al., 2018; Gollapudi & Panigrahi, 2019; Shin et al., 2023), caching (Narayanan et al., 2018; Rohatgi, 2020; Lykouris & Vassilvitskii, 2021; Im et al., 2022), scheduling (Gollapudi & Panigrahi, 2019; Lattanzi et al., 2020; Lassota et al., 2023), and matching (Antoniadis et al., 2023; Dinitz et al., 2021; Sakaue & Oki, 2022) by using predictions from oracles such as machine learning models.
-The main idea of these novel algorithms is to exploit the characteristics of the data by actively exploiting information not explicitly used by classical algorithms.
-In particular, our work is based on research about sorting with predictions.
-(Kraska et al., 2019) proposed a method for fast sorting that uses a model $\tilde{F}(q)$ trained to approximate the cumulative distribution function (CDF) $F(q)$.
-This first Learned Sort roughly sorts the array by placing each element $x$ of the array on the $nF(x)$-th entry of the output array.
-It then sorts the array completely by refining the roughly sorted array using insertion sort.
-This algorithm has a best case complexity of $\mathcal{O}(n)$, but a worst case complexity of $\mathcal{O}(n^{2})$ because of the exception handling for collisions and the last refining.
-Subsequently, implementations have been proposed that improve cache efficiency (Kristo et al., 2020) and are robust against key duplication (Kristo et al., 2021).
-These also use insertion sorting in the last refinement, resulting in a worst-case complexity of $\mathcal{O}(n^{2})$.
-The worst-case complexity can be reduced to $\mathcal{O}(n\log n)$ by using a sorting algorithm with small worst-case complexity, such as Introsort (Musser, 1997) or TimSort (McIlroy, 1993).
-However, this is the same complexity as many effective classical sorting algorithms, and thus does not explain the experimental success of Learned Sort.
-Some studies have focused on efficiently exploiting the predictions of the oracle, considering the oracle as a black box and not focusing on its inner details (Bai & Coester, 2023).
-This analysis does not take into account the computational complexity of training and inferring the oracle, so it cannot be directly applied to guarantee the overall computational complexity of sorting.
+Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section˜2.4).
+2.1 Algorithms with Machine Learning
+Our work lies in the emerging area of Learned Sort, where machine learning models accelerate sorting.
+The first approach by kraska2021sagedb trained a model $\tilde{F}(q)$ to approximate the CDF $F(q)$ and placed each element at position $n\tilde{F}(x)$, and then applies insertion sort to complete the ordering.
+While this yields $\mathcal{O}(n)$ in favorable cases, the insertion-based refinement causes a $\mathcal{O}(n^{2})$ worst case.
+Later implementations improved cache efficiency (kristo2020case), robustness to duplicates (kristo2021defeating), and even introduced cache-friendly CDF models tailored for sorting (ferragina2025balancedsort), but still relied on insertion sort and thus retained a quadratic worst-case complexity.
+Using stronger fallback algorithms such as Introsort (musser1997introspective) or TimSort (mcilroy1993optimistic) reduces this to $\mathcal{O}(n\log n)$, though this only matches classical comparison-based sorts and does not fully explain the empirical advantage of Learned Sort.
+Parallelization has also been explored: carvalho2022towards proposed IPLS, integrating learned partitioning into IPS4o (axtmann2017), and carvalho2023learnedsort framed LearnedSort as a sample sort with a parallel IPS4o implementation.
+These studies primarily highlight the engineering potential of learned sorting in parallel settings, which is complementary but orthogonal to our focus on providing theoretical guarantees for the sequential case.
+In a broader context, li2020strong showed that neural programs can achieve strong generalization and replicate efficient algorithmic behaviors such as sorting, demonstrating that machine learning can recover algorithmic complexity classes like $\mathcal{O}(n\log n)$ without explicit manual design.
+This provides a useful perspective on the generalizability of learned algorithms.
+Related to this, the field of algorithms with predictions studies how machine learning predictions accelerate classical algorithms (mitzenmacher2022algorithms), with applications in caching (narayanan2018deepcache; rohatgi2020near; lykouris2021competitive; im2022parsimonious), ski rental (purohit2018improving; gollapudi2019online; shin2023improved), scheduling (gollapudi2019online; lattanzi2020online; lassota2023minimalistic), and matching (antoniadis2023online; dinitz2021faster; sakaue2022discrete).
+Sorting with predictions has also been analyzed (lu2021generalized; chan2023generalized; erlebach2023sorting), including tight guarantees by bai2023sorting.
+In the context of algorithms with predictions, machine learning predictions are typically assumed to be available at no cost, and the models are treated as an opaque box.
+This exclusion contrasts with our problem setting, where we ensure that the computational complexity covers the entire process from receiving an unsorted array to returning a sorted array.
 2.2 Comparison Sorts
 Sorting algorithms that use comparisons between keys and require no other information about the keys are called comparison sorts.
-It can be proved that the worst complexity of a comparison sort is at least $\mathcal{O}(n\log n)$.
+It is well-known that the worst-case complexity of a comparison sort is at least $\Omega(n\log n)$.
 Commonly used comparison sorting algorithms include QuickSort, heap sort, merge sort, and insertion sort.
-The GNU Standard Template Library in C++ uses Introsort (Musser, 1997), an algorithm that combines quick sort, heap sort, and insertion sort.
-Java (Java, 2023) and Python up to version 3.10 (Peters, 2002) use TimSort (McIlroy, 1993), an improved version of merge sort.
-Python 3.11 and later use Powersort (Munro & Wild, 2018), a merge sort that determines a near-optimal merge order.
-Sample sort (Frazer & McKellar, 1970) is an extension of quick sort that uses multiple pivots, whereas quick sort uses only one pivot.
+The GNU Standard Template Library in C++ uses Introsort (musser1997introspective), an algorithm that combines QuickSort, heap sort, and insertion sort.
+Java (javalistsort2023) and Python up to version 3.10 (python_timsort2002) use TimSort (mcilroy1993optimistic), an improved version of merge sort.
+Python 3.11 and later use Powersort (munro2018nearly), a merge sort that determines a near-optimal merge order.
+Sample sort (frazer1970samplesort) extends QuickSort by using multiple pivots instead of one.
 Sample sort samples a small number of keys from the array, determines multiple pivots, and uses them to partition the array into multiple buckets.
 The partitioning is repeated recursively until the array is sufficiently small.
-The in-place parallel superscalar sample sort (Axtmann et al., 2022) is one of the most efficient sample sort implementation.
-They theoretically guarantee its computational and cache efficiency by proving a theorem about the probability that the pivots partition the array (nearly) equally.
-2.3 Distribution sorts
-Distribution sorts use information other than key comparison.
-Radix sort and counting sort are the most common types of distribution sorts.
+Among its implementations, the in-place parallel superscalar sample sort (axtmann2017) was introduced as a highly efficient parallel algorithm, and later engineering refinements further improved its performance (axtmann2022engineering).
+Its computational and cache efficiency is theoretically guaranteed by a theorem about the probability that the pivots partition the array (nearly) equally.
+2.3 Non-Comparison Sorts
+Non-comparison sorts use information other than key comparison.
+Radix sort and counting sort are the most common types of non-comparison sorts.
 Radix sort uses counting sort as a subroutine for each digit.
@@
 Thus, radix sort is particularly effective when the number of digits is small.
-There are several variants of radix sort, such as SkaSort (Skarupke, 2016), which is optimized using a technique called “American flag sort,” and RegionSort (Obeya et al., 2019), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
+There are several variants of radix sort, such as Spreadsort (ross2002spreadsort), which integrates the advantages of comparison sort into radix sort and is implemented in the Boost C++ Libraries, and RegionSort (obeya2019theoretically), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
+In addition, non-comparison sorting algorithms tailored for specific data types have been developed.
+For integer arrays, a deterministic algorithm with worst-case complexity of $\mathcal{O}(n\log\log n)$ (han2002deterministic) and a randomized algorithm with expected complexity of $\mathcal{O}(n\sqrt{\log\log n})$ (han2002integer) have been proposed.
+For real-valued arrays, recent advances have led to the development of a sorting algorithm with a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$ (han2020sorting).
+Our PCF Learned Sort also targets real-valued arrays and, under mild assumptions on the distribution, achieves an expected complexity of $\mathcal{O}(n\log\log n)$.
+Moreover, when the algorithm of han2020sorting is used as the internal sorting component, PCF Learned Sort has a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$, which matches that of han2020sorting.
 2.4 Learned Index
-Kraska et al.
-showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models, and named such novel data structures Learned Index (Kraska et al., 2018).
-Since then, various learning augmented B-trees (Wang et al., 2020; Kipf et al., 2020) and learning augmented Bloom filters (Mitzenmacher, 2018; Vaidya et al., 2021; Sato & Matsui, 2023) have been proposed.
-There are several researches on learning augmented B-trees whose performance is theoretically guaranteed.
-PGM-index (Ferragina & Vinciguerra, 2020) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$.
-(Zeighami & Shahabi, 2023) proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution.
+kraska2018case_li showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models and named such novel data structures Learned Index.
+Since then, learning augmented B-trees (wang2020sindex; kipf2020radixspline; ferragina2020pgm; zeighami2023distribution), Bloom filters (mitzenmacher2018model; vaidya2020partitioned; sato2023fast; sato2024fast), and even range-minimum query structures (ferragina2025fl) have been proposed.
+There are several works on learning augmented B-trees whose performance is theoretically guaranteed.
+PGM-index (ferragina2020pgm) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$.
+zeighami2023distribution proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution.
+More broadly, this line of research connects to earlier studies on compressed data structures with provable guarantees, such as the foundational work of sadakane2007compressed.
 3 Methods
+This section introduces our Learned Sort framework and its PCF-based instantiation.
+In Section˜3.1, we set up the notation, in Section˜3.2 we present the general framework, and in Section˜3.3 we instantiate it with a PCF-based CDF model to obtain concrete complexity guarantees.
+3.1 Notation and Setup
+Following the learned-sort paradigm kraska2021sagedb; kristo2020case; kristo2021defeating, our learned sort algorithm recursively applies model-based bucketing.
+Model-based bucketing partitions the array into multiple buckets using a CDF model trained on the input so that the bucket with the larger ID gets the larger value.
+Let $\mathcal{D}\subseteq\mathbb{R}$ be the domain of possible key values and $\bm{x}\in\mathcal{D}^{n}$ be the input array.
+For vector/array $\bm{v}$, we write $|\bm{v}|$ for its length (number of elements), e.g., $|\bm{x}|=n$.
+If $|\bm{x}|$ is smaller than a fixed threshold $\tau\in\mathbb{N}$, the algorithm falls back to a classic sorting algorithm.
+We refer to this fallback classical sorting algorithm as the standard sort.
+As the standard sort, we can use any sorting algorithm, such as IntroSort, QuickSort, or the algorithm of han2020sorting.
+For model-based bucketing, we define functions $\alpha(n),\beta(n),\gamma(n),\delta(n):\mathbb{N}\to\mathbb{N}$ as follows: $\alpha(n)$ is the number of samples used to train the CDF model, $\beta(n)$ is the number of PCF bins (a hyperparameter; see Section 3.3), $\gamma(n)+1$ is the number of buckets, and $\delta(n)$ is the threshold that determines the behavior after bucketing (see Section 3.2).
+For brevity, we write $\alpha,\beta,\gamma,\delta$.
+3.2 Learned-Sort Framework
 Figure 1: PCF Learned Sort: First, the input array is partitioned into $\gamma+1$ buckets using a CDF model-based method.
-Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a “standard” sorting algorithm of complexity $\mathcal{O}(n\log n)$ (e.g.
-IntroSort).
+Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a standard sort (e.g., IntroSort or QuickSort).
 Otherwise, the recursive model-based bucketing is repeated.
@@
 The CDF model used for bucketing is a Piecewise Constant Function (PCF).
-The function is constant within each interval and the interval widths are constant.
+The function is constant within each interval, and the interval widths are constant.
 Algorithm 1 The Learned Sort algorithm
+1:Input: $\bm{x}\in\mathbb{R}^{n}$Output: $\bm{x}_{\mathrm{sorted}}\in\mathbb{R}^{n}$ (the sorted version of $\bm{x}$)
+2:Subroutines:
 […]
 […]
+5:function Learned-Sort($\bm{x}$)
+6:  $n\leftarrow|\bm{x}|$
+7:  if $n<\tau$ then $\triangleright$ Small bucket
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-This section describes our PCF Learned Sort.
-First, in Section 3.1, we give an overview of the algorithm and the key theorems used to guarantee the expected and worst-case computational complexity.
-Next, in Section 3.2, we introduce the specific algorithm for PCF Learned Sort and give a guarantee on its computational complexity.
-3.1 Method Overview
-Our algorithm basically repeats recursive model-based bucketing until the array is small enough or the bucketing “fails.” In model-based bucketing, a CDF model is trained and then used to partition an array of length $n$ into approximately $\gamma$ buckets, satisfying the following two conditions: the elements in $i$-th bucket are smaller than any elements in $(i+1)$-th bucket, and the buckets have approximately the same size with high probability.
-If we set $\gamma=n^{c}$ (where $10<c<$ 1 is a constant) and the partitioning into buckets is done well, the size of each bucket will be approximately $n^{1-c}$, so the size of the bucket in the $k$-th recurrence depth will be approximately $n^{(1-c)^{k}}$.
-Here,
-$k\geq-\frac{\log\log n}{\log(1-c)}\Rightarrow n^{(1-c)^{k}}$ $\leq n^{(1-c)^{-\frac{\log\log n}{\log(1-c)}}}$ (1)
-$=n^{\frac{1}{\log n}}$ (2)
-$=e,$ (3)
-where $\log$ is the natural logarithm, so the base is $e$.
-That is, with a recursion depth of $k=\Omega(\log\log n)$, the size of bucket is small enough, i.e., the max recursion depth is $\mathcal{O}(\log\log n)$.
-Thus, if the expected computational complexity of the partition is $\mathcal{O}(n)$ and the probability that the bucketing “fails” is sufficiently small, the total computational complexity is $\mathcal{O}(n\log\log n)$.
-Under mild assumptions on the distribution, we can prove that this fast and accurate partition can be achieved by using the Piecewise Constant Function (PCF) as the CDF model.
-Let $\mathcal{D}~{}(\subset{\mathbb{R}})$ be the range of possible values of the input array and $\bm{x}~{}(\in\mathcal{D}^{n})$ be the input array.
-If the length of the input array, denoted as $|\bm{x}|$, is less than $\tau$, our algorithm sorts the input array using a “standard” sorting method with $\mathcal{O}(n\log n)$ complexity, such as IntroSort, where $\tau$ is a predetermined constant.
-Otherwise, if $|\bm{x}|\geq\tau$, model-based bucketing is performed.
+9:  $\tilde{F}(q)\leftarrow\textsc{CDF-Model}(\bm{x})$ $\triangleright$ Model-based bucketing
+10:  $\bm{c}_{1}\leftarrow[~],~\bm{c}_{2}\leftarrow[~],~\dots~,~\bm{c}_{\gamma+1}\leftarrow[~]$
+11:  for $i=1,2,\dots,n$
+12:   $j\leftarrow\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1$
+13:   $\bm{c}_{j}.\textsc{Append}(x_{i})$
+14:  for $j=1,2,\dots,\gamma+1$ $\triangleright$ Recursively sort and concatenate
+15:   if $|\bm{c}_{j}|\geq\delta$ then
+16:     $\bm{c}_{j}\leftarrow\textsc{Standard-Sort}(\bm{c}_{j})$
+17:   else
+18:     $\bm{c}_{j}\leftarrow\textsc{Learned-Sort}(\bm{c}_{j})$
+19:  $\bm{x}_{\mathrm{sorted}}\leftarrow\textsc{Concatenate}(\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{\gamma+1})$
+20:  return $\bm{x}_{\mathrm{sorted}}$
+Algorithm description.
+At a high level, our algorithm is a recursive model-based bucketing scheme augmented with an exception-handling mechanism for bucketing “failures.” This exception-handling mechanism greatly simplifies the analysis of both expected and worst-case complexity.
+The overall workflow is visualized in Figure˜1, and the pseudocode is given in Algorithm˜1.
+If the length of the input array is less than $\tau$, our algorithm sorts the input array using a standard sort (e.g., IntroSort or QuickSort). Otherwise, model-based bucketing is performed.
 The model-based bucketing method $\mathcal{M}$ takes an input array $\bm{x}$ and partitions it into several buckets.
-First, it determines the parameter $\gamma~{}(\in\mathbb{N})$, which determines the number of buckets.
-$\gamma$ is determined as a function of the array length $n$.
-The number of buckets returned by $\mathcal{M}$ is $\gamma+1$.
-Next, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\mapsto[0,1]$.
-The $\tilde{F}(q)$ is trained to predict the empirical cumulative distribution function of $\bm{x}$ defined as:
-$F_{\bm{x}}(q)\coloneqq\frac{|\{i\in\{1,\dots,n\}\mid x_{i}\leq q\}|}{|\bm{x}|}.$ (4)
-The specific model and training method is explained in Section 3.2.
-Finally, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets.
+First, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\to[0,1]$.
+The $\tilde{F}(q)$ is trained to approximate the empirical CDF $F_{\bm{x}}(q)=|{i\in\{1,\dots,n\}\mid x_{i}\leq q}|/|\bm{x}|$.
+Any non-decreasing model can serve as the CDF model $\tilde{F}(q)$ (e.g., linear models, monotonic MLPs, or the PCF introduced in Section˜3.3).
+After training, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets.
 All $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$, are initialized to be empty, and then for each $i\in\{1,\dots,n\}$, $x_{i}$ is appended to $\bm{c}_{\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1}$.
-This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., |$|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
-The CDF model $\tilde{F}$ must be a non-decreasing function to guarantee the computational complexity of our algorithm. In this case, the bucket with the larger ID gets the larger value, i.e.,
-$p\in\bm{c}_{j},q\in\bm{c}_{k},1\leq j<k\leq\gamma+1\Rightarrow p<q.$ (5)
+This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., $|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
+We restrict the CDF model $\tilde{F}$ to non-decreasing functions to ensure that the bucket with the larger ID gets the larger value, i.e., $p\in\bm{c}_{j}\;\land\;q\in\bm{c}_{k}\;\land\;j<k\;\;\Rightarrow\;\;p<q$.
 This means that each bucket is responsible for a disjoint and continuous interval.
 Let $t_{j}=\min_{x\in\bm{c}_{j}}x~(j=1,\dots,\gamma+1),~t_{\gamma+2}=\infty$, then the $j$-th bucket $\bm{c}_{j}$ ($j=1,\dots,\gamma+1$) is responsible for a continuous interval $\mathcal{I}_{j}\coloneqq[t_{j},t_{j+1})$.
-After model-based bucketing, our algorithm determines for each bucket whether the bucketing was “successful” or not.
-For each $j=1,\dots,\gamma+1$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$, where $\delta$ is an integer determined by $n$.
-If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the “standard” sorting method (e.g., IntroSort).
+After model-based bucketing, our algorithm determines for each bucket whether the bucketing “succeeds” or “fails.” For each $j\in\{1,\dots,\gamma+1\}$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$.
+If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the standard sort (e.g., IntroSort or QuickSort).
 If $|\bm{c}_{j}|<\delta$ (which means the bucketing “succeeds”), the bucket is sorted by recursively calling our Learned Sort algorithm.
-Note that the parameters such as $\gamma$ and $\delta$ are redetermined according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket.
-After each bucket is sorted, the buckets are concatenated.
-The input array $\bm{x}$ is sorted by the above procedure.
-The basic idea of our algorithm is visualized in Figure 1, and the pseudocode is given in Algorithm 1.
-The following is a lemma about the worst-case complexity of our Learnd Sort as defined above.
+Note that the parameters such as $\gamma$ and $\delta$ are redetermined for each recursion according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket.
+Finally, the sorted buckets are concatenated to form the sorted array.
+Worst-case complexity.
+The following lemma guarantees the worst-case complexity of our Learned Sort.
 Lemma 3.1.
-Assume that there exists an model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$.
-Also, assume that the “standard” sort algorithm used for the too large or too small buckets has a worst-case complexity of $\mathcal{O}(n\log n)$.
-The worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $10<d<$ 1) is $\mathcal{O}(n\log n)$.
-This lemma can be intuitively shown from the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) each element of the input array $\bm{x}$ undergoes several bucketing and only one “standard” sort.
-(i) can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$, and (ii) is evident from the algorithm’s design since the buckets sorted by “standard” sort are now left only to be concatenated.
-The exact proof is given in Section A.1.
-Next, we introduce an important lemma about the expected complexity of our Learned Sort. The following assumption is necessary to guarantee the expected complexity.
+Assume that there exists a model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$.
+Also, assume that the standard sort has a worst-case complexity of $\mathcal{O}(nU(n))$, where $U(n)$ is a non-decreasing function.
+Then, the worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $0<d<1$) is $\mathcal{O}(nU(n)+n\log\log n)$.
+This lemma can be intuitively shown from the following two points: (1) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (2) each element of the input array $\bm{x}$ undergoes several bucketing and only one standard sort.
+The first point can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$.
+The second point is evident from the algorithm’s design since the buckets sorted by standard sort are now left only to be concatenated.
+The exact proof is given in Section˜A.1.
+Note that this guarantee critically relies on the exception-handling mechanism: without it, the recursion depth could reach $\Theta(n)$ and the worst-case time would degrade to $\Theta(n^{2})$, whereas with it the recursion depth is always bounded by $O(\log\log n)$.
+Expected complexity.
+Next, we introduce a lemma about the expected computational complexity of our Learned Sort. The following assumption is necessary to guarantee the expected computational complexity.
 Assumption 3.2.
-The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\mapsto\mathbb{R}_{\geq 0}$.
-We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\mapsto\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subset\mathcal{D}$, i.e.,
-[…]
+The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\to\mathbb{R}_{\geq 0}$.
+We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\to\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subseteq\mathcal{D}$, i.e., $f_{\mathcal{I}}(x)\coloneqq\frac{f(x)}{\int_{\mathcal{I}}f(y)dy}$.
 The expected computational complexity of our proposed Learned Sort is guaranteed by the following lemma.
@@
 Let $\bm{x}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ be the array formed by sampling $n$ times independently according to $f_{\mathcal{I}}(x)$.
-Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~{}(\in(0,1))$ that satisfy the following for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
-  • $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$, with an expected complexity of $\mathcal{O}(n)$.
-That is, $\mathcal{M}$ can take the array $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ as input and divide it into $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{j=\gamma+1}$ satisfying Equation 5, with an expected complexity of $\mathcal{O}(n)$.
-  • $\gamma+1=\mathcal{O}(n)$.
-  • $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$.
-The expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
-This lemma can be proved intuitively by the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$.
-(i) is the same as in the explanation of the proof of Lemma 3.1.
-(ii) can be shown from the fact that the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases.
-See Section A.2 for the exact proof.
-Note that the assumption of Lemma 3.3 includes “$\mathcal{M}$ works well with high probability for any $\mathcal{I}$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing.
-The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~{}(\subset\mathcal{D})$.
+Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~(\in(0,1))$ that satisfy the following three conditions for any interval $\mathcal{I}~(\subseteq\mathcal{D})$: (i) $\mathcal{M}$ can perform bucketing (including model training and inferences) on an array of length $n$, with an expected complexity of $\mathcal{O}(n)$, (ii) $\gamma+1=\mathcal{O}(n)$, and (iii) $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$.
+Also, assume that the standard sort has an expected complexity of $\mathcal{O}(n\log n)$.
+Then, the expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
+This lemma can be proved intuitively by the following two points: (1) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (2) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$.
+The first point is the same as in the explanation of the proof of Lemma˜3.1.
+The second point holds because the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases.
+See Section˜A.2 for the exact proof.
+Note that the assumption of Lemma˜3.3 includes $\mathcal{M}$ works well with high probability for any $\mathcal{I}~(\subseteq\mathcal{D})$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing.
+The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~(\subseteq\mathcal{D})$.
 3.3 PCF Learned Sort
-We propose PCF Learned Sort as an implementation that satisfies the assumptions of Lemma 3.1 and Lemma 3.3, and thus has a guarantee that the worst-case complexity is $\mathcal{O}(n\log n)$ and the expected complexity is $\mathcal{O}(n\log\log n)$.
-PCF Learned Sort approximates the CDF by a Piecewise Constant Function (PCF).
-Piecewise constant function is a function that has intervals of equal width and outputs the same value in each interval (the right side of Figure 1).
-The study that develops a Learned Index with a theoretical guarantee on its complexity (Zeighami & Shahabi, 2023) also used PCF as a CDF model.
-The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ in the following way.
-First, the parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, the length of the input array.
-The parameter $\alpha$ is the number of samples to train the CDF model, and $\beta$ is the number of intervals in the PCF.
-Next, the piecewise constant function is trained by counting the number of samples in each interval.
-Using $x_{\mathrm{min}}=\min_{i}{x_{i}}$ and $x_{\mathrm{max}}=\max_{i}{x_{i}}$, let $i(x)$ be a function defined as follows:
-$i(x)=\left\lfloor\frac{x-x_{\mathrm{min}}}{x_{\mathrm{max}}-x_{\mathrm{min}}}\beta\right\rfloor+1.$ (7)
-$\alpha$ samples are taken at random from $\bm{x}$ to form $\bm{a}~{}(\in\mathcal{D}^{\alpha})$, and then $i(x)$ is used to form the array $\bm{b}~{}(\in{\mathbb{Z}_{\geq 0}}^{\beta+1})$ defined as follows:
-$b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|.$ (8)
-This counting is corresponding to the training of the PCF.
+We now instantiate the framework with PCF Learned Sort, which satisfies the assumptions of Lemma˜3.1 and Lemma˜3.3, thereby providing both worst-case and expected-time guarantees.
+PCF Learned Sort approximates the CDF using a Piecewise Constant Function (PCF) with $\beta$ equal-width bins; the model output is constant within each bin (see the right panel of Figure 1).
+The study that develops a Learned Index with a theoretical guarantee on its complexity (zeighami2023distribution) also used PCF as a CDF model.
+While our framework admits more expressive CDF models (such as spline-based or neural models), we focus on PCF due to its minimal training and inference cost and the tractability of its theoretical analysis.
+The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ as follows.
+The parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, where $\alpha$ is the number of samples used to train the model and $\beta$ is the number of intervals in the PCF.
+The PCF is trained by counting the number of samples in each interval.
+We define $i(x)=\lfloor(x-x_{\min})\beta/(x_{\max}-x_{\min})\rfloor+1$, where $x_{\min}=\min_{i}x_{i}$ and $x_{\max}=\max_{i}x_{i}$.
+From $\bm{x}$, $\alpha$ samples are taken to form $\bm{a}\in\mathcal{D}^{\alpha}$, and $i(x)$ is used to construct $\bm{b}\in\mathbb{Z}_{\geq 0}^{\beta+1}$ with $b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|$.
+This counting procedure trains the PCF.
 Note that $\bm{b}$ is an non-decreasing non-negative array and $b_{\beta+1}=\alpha$, i.e., $0\leq b_{1}\leq b_{2}\leq\dots\leq b_{\beta+1}=\alpha$.
-The inference for the CDF model $\tilde{F}(x)$ is performed using $i(x)$, $\bm{b}$, and the following equation:
-$\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}.$ (9)
-Since $i(x)$ is a non-decreasing function and $\bm{b}$ is also a non-decreasing array, $\tilde{F}(x)$ is a non-decreasing function.
+Inference for $\tilde{F}(x)$ is then given by $\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}$.
+Since $i(x)$ and $\bm{b}$ are non-decreasing, $\tilde{F}(x)$ is also non-decreasing.
 Also, $0\leq\tilde{F}(x)\leq 1$ because $0\leq b_{i}\leq\alpha$ for every $i$.
@@
 That is, $x\in\mathcal{D}\;\;\Rightarrow\;\;\sigma_{1}\leq f(x)\leq\sigma_{2}$.
-Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
-$\begin{split}&K\geq 1\\ \Rightarrow&\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},\end{split}$ (10)
-where
-$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}.$ (11)
+Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~(\subseteq\mathcal{D})$:
+$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}\geq 1\;\;\Rightarrow\;\;\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (1)
 The proof of this lemma is based on and combines proofs from two existing studies.
 The first is Lemma 5.2.
-from a study of IPS4(Axtmann et al., 2022), an efficient sample sort.
+from a study of IPS4(axtmann2022engineering), an efficient sample sort.
 This lemma guarantees the probability of a “successful recursion step” when selecting pivots from samples and using them to perform a partition.
-Although the definition of “successful recursion step” in this paper is similar to our definition of “successful bucketing” (more precisely, in the special case of $\delta=3n/\gamma$), this lemma is for the method that do not use the CDF model, so the proof cannot be applied directly to our case.
+This lemma is for the method that does not use the CDF model, so the proof cannot be applied directly to our case.
 Another proof we refer to is the proof of Lemma 4.5.
-from a study that addressed the computational complexity guarantee of the Learned Index (Zeighami & Shahabi, 2023).
+from a study that addressed the computational complexity guarantee of the Learned Index (zeighami2023distribution).
 This lemma provides a probabilistic guarantee for the error between the output of the PCF and the empirical CDF.
 Some modifications are required to adapt it to the context of sorting and to attribute it to the probability of bucketing failure, i.e., $\Pr[\exists j,|\bm{c}_{j}|>\delta]$.
-By appropriately combining the proofs of these two lemmas, Lemma 3.4 can be proved.
-The exact proof is given in Section A.3.
-Using Lemma 3.1, Lemma 3.3 and Lemma 3.4, we can prove the following theorems.
+By appropriately combining the proofs of these two lemmas, Lemma˜3.4 is proved.
+The exact proof is given in Section˜A.3.
+Here, we emphasize that the assumption of this lemma, $0<\sigma_{1}\leq\sigma_{2}<\infty$, is sufficiently reasonable and “mild” as described in (zeighami2023distribution).
+It asserts that the probability density function $f(x)$ is both bounded and nonzero over its domain $\mathcal{D}$.
+This class of distributions covers the majority of real-world scenarios because real-world data is commonly derived from bounded and continuous phenomena, e.g., age, grades, and data over a period of time.
+Empirically, zeighami2023distribution further suggest that the ratio $\sigma_{2}/\sigma_{1}$ tends to remain close to 1 in a wide variety of practical datasets, and even for more challenging cases like OSM, the ratio still appears to remain at most around 20.
+Using Lemma˜3.1, Lemma˜3.3, and Lemma˜3.4, we can prove the following theorems.
 Theorem 3.5.
-If we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
+If $\mathcal{M}_{\mathrm{PCF}}$ is the bucketing method, the worst-case complexity of standard sort is $\mathcal{O}(nU(n))$ (where $U(n)$ is a non-decreasing function), and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the worst-case complexity of PCF Learned Sort is $\mathcal{O}(nU(n)+n\log\log n)$.
 Proof.
-When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$.
-This is because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element.
-Therefore, from Lemma 3.1, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
+When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$ because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element.
+Therefore, since $\gamma+1=\mathcal{O}(n)$, the worst-case complexity of standard sort is $\mathcal{O}(nU(n))$, and $\delta=\lfloor n^{3/4}\rfloor$, we can prove the worst-case complexity of PCF Learned Sort is $\mathcal{O}(nU(n)+n\log\log n)$ by Lemma˜3.1.
 ∎
@@
 Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
-Then, if we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
+Then, if $\mathcal{M}_{\mathrm{PCF}}$ is the bucketing method, the expected complexity of standard sort is $\mathcal{O}(n\log n)$, and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
 Proof.
@@
 Since $K=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K\geq 1$ for sufficiently large $n$, and
-$\begin{split}\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}&=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\\ &\leq\mathcal{O}\left(\frac{1}{\log n}\right).\end{split}$ (12)
-Therefore, from Lemma 3.3, the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
-Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort, since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$ and $\delta$, are determined without the knowledge.
+$\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\leq\mathcal{O}\left(\frac{1}{\log n}\right).$ (2)
+Given that $\gamma+1=\mathcal{O}(n)$, the expected complexity of standard sort is $\mathcal{O}(n\log n)$, and $\delta=\lfloor n^{3/4}\rfloor$, it follows from Lemma˜3.3 and Lemma˜3.4 that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
+∎
+Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$, and $\delta$, are determined without any prior knowledge.
 In other words, PCF Learned Sort can sort in expected $\mathcal{O}(n\log\log n)$ complexity as long as $0<\sigma_{1}\leq f(x)\leq\sigma_{2}<\infty$, even if it does not know the exact value of $\sigma_{1}$ and $\sigma_{2}$.
-If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem 3.6, i.e., $\sigma_{1}=0$ or $\sigma_{2}=\infty$, then the expected complexity of PCF Learned Sort is increased to $\mathcal{O}(n\log n)$, but from Theorem 3.5, it cannot be greater than $\mathcal{O}(n\log n)$.
+If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem˜3.6, the expected complexity of PCF Learned Sort increases to $\mathcal{O}(nU(n)+n\log\log n)$.
+Such scenarios arise, for instance, in heavy-tailed distributions, datasets with extremely sparse regions ($\sigma_{1}=0$), or cases where the density is highly concentrated at particular values ($\sigma_{2}=\infty$).
+However, thanks to the worst-case bound in Theorem˜3.5, the complexity never exceeds $\mathcal{O}(nU(n)+n\log\log n)$.
+Thus, the algorithm remains robust even under distributions that do not fully satisfy the assumption.
+Alternative CDF Models.
+Our framework is not limited to PCF; it can also accommodate more expressive CDF models.
+For instance, we can replace PCF with a spline-based model that approximates the CDF by interpolating the empirical distribution at bin boundaries.
+Concretely, given $\beta$ intervals, we evaluate the empirical CDF at the endpoints of each interval and then construct a piecewise linear spline that connects these values, yielding a continuous and non-decreasing CDF approximation.
+We refer to the algorithm that applies this spline-based CDF model within our Learned Sort framework as Spline Learned Sort.
+We show that spline-based models constructed in this way still satisfy the assumptions of Lemma˜3.3, and thus the expected time complexity guarantee of $\mathcal{O}(n\log\log n)$ remains valid.
+A detailed theoretical proof for the spline-based case is provided in Section˜A.5.
 4 Experiments
-In this section, we confirm our theorems experimentally.
-First, we confirm that the computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real data.
-Then, we conduct experiments with various parameter settings and experimentally confirm Lemma 3.4, a lemma that bounds the probability of bucketing failure and is an important lemma to guarantee the expected computational complexity of PCF Learned Sort.
-4.1 Results on Synthetic Datasets
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Figure 2: Number of operations to sort the synthetic array. Below each graph is a histogram that visualizes the distribution of each dataset.
-We experimented with datasets created from the following four types of synthetic distributions: uniform distribution ($\mathsf{min}=0,\mathsf{max}=1$), normal distribution ($\mu=0,\sigma=1$), exponential distribution ($\lambda=1$), lognormal distribution ($\mu=0,\sigma=1$).
-We took $n=10^{3},\dots,10^{8}$ samples from each distribution independently, and meticulously counted the total number of basic operations to sort the sampled array.
-Here, the basic operations consists of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access.
-We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware (this is the same idea as the metric selection in the experiment of  (Zeighami & Shahabi, 2023)).
-The parameters of PCF Learned Sort are set as in Theorem 3.5 and Theorem 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$.
-As the “standard” sort algorithm used in PCF Learned Sort, we used quick sort, which is known to have an expected computational complexity of $\mathcal{O}(n\log n)$.
-We then compared the number of operations between PCF Learned Sort and the (plain) quick sort.
-Figure 2 shows the number of operations divided by the length of input array, $n$.
-It shows the mean and standard deviation of the $10$ measurements for each condition.
-We see that our PCF Learned Sort has up to 2.8 times fewer operations than quick sort.
-Also, while the graph of quick sort is almost linear, the graph of PCF Learned Sort is almost flat.
-This suggests that PCF Learned Sort has a computational complexity much smaller than $\mathcal{O}(n\log n)$.
-4.2 Results on Real Datasets
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Refer to captionRefer to caption
-Figure 3: Number of operations to sort the real data array. Below each graph is a histogram that visualizes the distribution of each dataset.
-We used four real datasets, NYC, Wiki, OSM, and Books, which are described below.
-NYC: pick-up datetimes in the yellow taxi trip records (Kristo, 2021).
-Wiki: Wikipedia article edit timestamps (Marcus et al., 2020).
-OSM: uniformly sampled OpenStreetMap locations represented as Google S2 CellIds (Marcus et al., 2020).
-Books: book sale popularity data from Amazon (Marcus et al., 2020).
-For each dataset, we randomly sample $n=10^{3},\dots,10^{8}$ elements, shuffle the sampled elements, and then use it as an input array to examine the performance of the algorithms.
-As in the case of the synthetic dataset, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and quick sort is used as “standard” sort.
-Figure 3 shows plots of the number of operations divided by the length of input array, $n$.
-Again, the mean and standard deviation of $10$ measurements are shown.
-We see that our PCF Learned Sort has up to 2.7 times fewer operations than quick sort (as in the case of the synthetic dataset), and that PCF Learned Sort has much less computational complexity than $\mathcal{O}(n\log n)$.
-We can also see that the number of operations required for the OSM dataset is about $1.25$ times larger than that for the other datasets.
-Although not mentioned explicitly, the same tendency is observed in the experiments of existing Learned Sort studies (Kristo et al., 2020, 2021).
-Our theoretical analysis can explain this phenomenon from the perspective of $\sigma_{2}/\sigma_{1}$, in other words, the strength of the distributional bias.
-From Equation 11 and Equation 10, the larger $\sigma_{2}/\sigma_{1}$ is, the smaller $K$ is, and the larger (the upper bound of) the probability of bucketing failure.
-Therefore, the larger $\sigma_{2}/\sigma_{1}$ is, the more the computational complexity will move away from $\mathcal{O}(n\log\log n)$ and closer to the worst-case computational complexity, $\mathcal{O}(n\log n)$.
-In fact, the histograms of each dataset show that the OSM dataset has sharp peaks and many intervals with almost no data, indicating that it has a distribution with large $\sigma_{2}/\sigma_{1}$.
-4.3 Confirmation of Lemma 3.4
-Figure 4: Heatmap showing the experimental frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$.
+In this section, we empirically validate our theoretical results.
+First, in Section˜4.1, we confirm that the complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real datasets.
+Then, in Section˜4.2, we conduct experiments with various parameter settings and empirically confirm Lemma˜3.4, a lemma that bounds the probability of bucketing failure and plays a crucial role in guaranteeing the expected complexity of PCF Learned Sort.
+Finally, in Section˜4.3, we present from sorting time measurements.
+Setup.
+We experimented with synthetic datasets created from the following four distributions: uniform ($\mathsf{min}=0,\mathsf{max}=1$), normal ($\mu=0,\sigma=1$), exponential ($\lambda=1$), lognormal ($\mu=0,\sigma=1$).
+The input array was generated by independently taking $n$ samples from each distribution.
+Only the uniform distribution satisfies the theoretical assumptions required for our complexity guarantees.
+The other distributions violate these assumptions, but we include them to evaluate the empirical robustness of PCF Learned Sort.
+We also used the following 16 real datasets, including Chicago [Start, Tot] (chic), NYC [Pickup, Dist, Tot] (nyc), SOF [Humidity, Pressure, Temperature] (sof), Wiki, OSM, Books, Face (sosd-vldb), and Stocks [Volume, Open, Date, Low] (stks).
+Further dataset details are provided in Appendix B.
+For each dataset, we randomly sample $n$ elements, shuffle them, and use them as an input array to examine the complexity of the sort algorithms.
+Since these are real-world datasets, we cannot definitively determine whether they satisfy our theoretical assumptions.
+However, as shown in the histograms in Figure˜2, Chicago [Start], NYC [Pickup, Tot], SOF [Humidity], and Face appear to distribute values across a relatively dense and continuous domain, aligning well with our assumptions.
+In contrast, other datasets exhibit long tails or sparse regions, suggesting that our assumptions may not hold.
+Furthermore, note that for several datasets (Chicago [Start, Tot], NYC [Dist, Tot], SOF [Humidity, Pressure, Temperature], and Stocks [Volume, Open, Date, Low]), the number of unique values is extremely small (less than 3.2% of the total elements, as detailed in Appendix B).
+All experiments were run on a Linux machine equipped with an Intel®  Core™  i9-11900H CPU @ 2.50GHz and 62GB of memory. GCC version 9.4.0 was used for compilation, employing the -O3 optimization flag.
+4.1 Computational Complexity of PCF Learned Sort
+We meticulously counted the total number of basic operations for sorting the input array to observe the computational complexity of each sorting algorithm.
+Here, the basic operations consist of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access.
+We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware.
+This is the same idea as the metric selection in the experiment of  (zeighami2023distribution).
+The parameters of PCF Learned Sort are set as in Theorems˜3.5 and 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and $\tau=100$.
+As the standard sort, we used QuickSort, which has an expected complexity of $\mathcal{O}(n\log n)$.
+For ease of implementation and straightforward measurement of comparison counts, we chose QuickSort.
+We compared our PCF Learned Sort against (plain) QuickSort, radix sort, and Learned Sort 2.0 (kristo2021defeating-github).
+Learned Sort 2.0 was selected as the learned-method baseline because it allows relatively simple counting of operations (a broader comparison appears in the sorting time experiments of Section˜4.3).
+Figure 2: Number of operations to sort the array.
+Below each graph is a histogram visualizing the distribution of each dataset.
+The standard deviation of the $10$ measurements is represented by the shaded area.
+Our PCF Learned Sort consistently achieves a complexity lower than $\mathcal{O}(n\log n)$, while Learned Sort 2.0 (kristo2021defeating), which has $\mathcal{O}(n^{2})$ worst-case complexity, occasionally requires huge operations.
+Figure˜2 shows the number of operations divided by the length of the input array, $n$.
+Each point represents the average over 10 runs, with the shaded region indicating the standard deviation.
+Note that the horizontal axis is logarithmic.
+As a result, the curve for QuickSort (with $\mathcal{O}(n\log n)$ complexity) appears approximately linear in this plot for synthetic datasets and real-world datasets with few duplicates (i.e., NYC [Pickup], Wiki, OSM, Books, and Face).
+In contrast, the curve for PCF Learned Sort is nearly flat, suggesting a complexity significantly lower than $\mathcal{O}(n\log n)$.
+PCF Learned Sort consistently requires the fewest operations across all conditions, achieving up to 2.8 times fewer operations than QuickSort.
+These results confirm not only our theoretical analysis but also the robustness of PCF Learned Sort under assumption-violating scenarios.
+For datasets with many duplicates, QuickSort, Learned Sort 2.0 (kristo2021defeating), and PCF Learned Sort all exhibit relatively few operations.
+This behavior arises because highly duplicated datasets often result in buckets containing only one distinct value, allowing the algorithms to terminate early.
+Even under such conditions, PCF Learned Sort performs fewer operations than the other algorithms.
+The curve for radix sort is also flat but lies consistently above that of PCF Learned Sort.
+This is due to the difference in partitioning strategies.
+While the radix sort partitions at a fixed granularity, our PCF Learned Sort performs partitioning using intervals that are adaptively set by the learning model.
+In particular, on the SOF [Temperature] dataset with medium input sizes ($2\times 10^{5}<n\leq 2\times 10^{7}$), Learned Sort 2.0 incurs extremely high costs: while all other methods keep the number of operations per $n$ below 100, it can exceed 50,00050{,}000.
+This shows that Learned Sort methods without worst-case guarantees may suffer from pathological overhead.
+By contrast, our method maintains both expected and worst-case guarantees, ensuring robust performance across diverse datasets.
+4.2 Confirmation of Lemma˜3.4
+Figure 3: Heatmap showing the empirical frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$.
 The variables $a,b,c,d$, except those on the x- and y-axes, were set to $0.75$.
-The white dotted line represents the parameters that make the right side of Equation 10 equal to 0.5.
-Lemma 3.4 bounds the probability that a bucket of size greater than $\delta$ exists.
+The white dotted line represents the parameters that make the right side of Equation˜1 equal to 0.5.
+The close alignment between this white dotted line and the actual success/fail boundery suggests the theoretical bound by is Equation˜1 reasonably tight.
+Lemma˜3.4 bounds the probability that a bucket of size greater than $\delta$ exists.
 This is an important lemma that allows us to guarantee the expected computational complexity of PCF Learned Sort.
-Here, we experimentally confirm that this upper bound is appropriate.
+Here, we empirically confirm that this upper bound is appropriate.
 We have experimented with $\alpha=\lfloor n^{a}\rfloor,\beta=\lfloor n^{b}\rfloor,\gamma=\lfloor n^{c}\rfloor,\delta=\lfloor n^{d}\rfloor$, varying $a,b,c,d$ from 0.05 to 0.95 at 0.05 intervals.
 For each $a,b,c,d$ setting, the following was repeated $100$ times: we took $n=10^{6}$ elements from the uniform distribution to form the input array and divided the array into $\gamma+1$ buckets by $\mathcal{M}_{\mathrm{PCF}}$, and checked whether or not $\exists j,|\bm{c}_{j}|>\delta$.
-Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the experimental frequency at which bucketing “fails.”
-Heat maps in Figure 4 show the experimental frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed and the other two parameters are varied.
-The value of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure 4 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$.
-The white dotted line represents the parameter so that the right side of Equation 10 is 0.5.
-That is, Lemma 3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
-We can see that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often.
-In other words, we can see that the theoretical upper bound from Lemma 3.4 agrees well (to some extent) with the actual probability.
-We can also confirm that, as Lemma 3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
-5 Limitation and Future Work
-We proposed PCF Learned Sort, whose expected computational complexity can be proved to be $\mathcal{O}(n\log\log n)$ under the assumption that there exist $0<\sigma_{1}$ and $\sigma_{2}<\infty$ satisfying $\sigma_{1}<f(x)<\sigma_{2}$.
-Thus, our proof is not applicable to the case of a distribution where there exists $x\in\mathcal{D}$ such that $f(x)=0$ or $f(x)=\infty$.
-The development of a theory and/or algorithm that is applicable for a wider class of distributions is a future work.
-Other future work may include incorporating a more advanced CDF approximation method with a theoretical guarantee into our Learned Sort algorithm.
-Using a different CDF model and a bucketing algorithm that satisfies the assumptions in Lemma 3.3, we may be able to develop an even more efficient sorting algorithm.
-In other words, a more accurate and faster CDF approximation method can lead to an improvement of the Learned Sort.
+Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the empirical frequency at which bucketing “fails.”
+Heat maps in Figure˜3 show the empirical frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed, and the other two parameters are varied.
+The values of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure˜3 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$.
+The white dotted line represents the parameter so that the right side of Equation˜1 is 0.5.
+That is, Lemma˜3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
+We observe that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often.
+This suggests that the theoretical upper bound from Lemma˜3.4 agrees well (to some extent) with the actual probability.
+We can also confirm that, as Lemma˜3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
+4.3 Experiments on Sorting Time
+We empirically compare the sorting time of our PCF Learned Sort against several baselines.
+Note that the metric used in Section˜4.1, the number of operations, does not change depending on the machine or compilation method, but the sorting time does.
+As the standard sort used in PCF Learned Sort, we used std::sort, which has a worst-case complexity of $\mathcal{O}(n\log n)$.
+We compared our PCF Learned Sort with std::sort, radix sort, boost::sort::spreadsort::float_sort (Boost C++ implementation of Spreadsort (ross2002spreadsort)), and Learned Sort 2.0 (kristo2021defeating-github).
+In addition, we evaluated more recent state-of-the-art learned sorting algorithms (ferragina2025balancedsort)—Balanced Learned Sort (BLS), Unbalanced Learned Sort (ULS), and Learned Sort 2.1—as well as IS4o (axtmann2017), one of the state-of-the-art non-learned sequential sample sort algorithms.
+Figure 4: Time to sort the array.
+The standard deviation of the $10$ measurements is represented by the shaded area.
+Our PCF Learned Sort is significantly faster than std::sort, and more importantly, it maintains robust performance across all datasets, whereas other learned sorts without worst-case guarantees can suffer catastrophic slowdowns, as seen with Learned Sort 2.0 on the SOF [Temperature] dataset.
+Figure˜4 shows the sorting time divided by $n$.
+It shows the mean and standard deviation of the $10$ measurements for each condition.
+Note that the horizontal axis is logarithmic, and therefore the curve of std::sort, which has a complexity of $\mathcal{O}(n\log n)$, is almost linear in this plot for synthetic datasets and real-world datasets with few duplicates.
+In contrast, our PCF Learned Sort graph shows a relatively slow increase, suggesting that PCF Learned Sort has a complexity much smaller than $\mathcal{O}(n\log n)$.
+We see that our PCF Learned Sort is up to 2.5 times faster than std::sort.
+Furthermore, we find that for relatively large $n$ ($10^{6}$), our PCF Learned Sort usually outperforms not only radix sort but also Spreadsort, an algorithm that cleverly incorporates the advantage of comparison sort into radix sort.
+The figure also shows that highly optimized methods like IS4o and other learned sorts often outperform PCF Learned Sort in average speed.
+This is because these methods are highly optimized implementations that consider factors like cache efficiency, whereas our implementation prioritizes providing rigorous theoretical guarantees.
+However, this speed comes at the cost of robustness: learned sorts without worst-case guarantees can suffer catastrophic slowdowns.
+For example, on the SOF [Temperature] dataset with $n=2\times 10^{7}$, Learned Sort 2.0 took up to 326.4 seconds, while our method consistently finished in 0.45 seconds.
+Moreover, as shown in our adversarial analysis (Appendix˜C), other recent learned sorts (BLS, ULS, Learned Sort 2.1) also exhibit vulnerabilities.
+These results highlight the practical value of our theoretical guarantees.
+To better understand runtime behavior, we profiled PCF Learned Sort across its stages (Appendix˜D).
+The results show that while the training stage is consistently lightweight across datasets and input sizes, the bucketing stage often becomes the bottleneck.
+The dominant cost varies depending on both $n$ and the dataset characteristics, suggesting opportunities for further optimization.
+5 Discussion
+Comparison with (zeighami2024theoretical).
+Most recently, a concurrent work (zeighami2024theoretical) introduced a theoretical framework for learned database operations, including sorting, indexing, and cardinality estimation.
+A key strength of their approach is the formal definition of “distribution learnability” and its applicability to a broad class of distributions, including those subject to distribution shifts.
+Using this framework, they developed a Learned Sort algorithm with an expected running time of $O(n\log\log n)$ under certain distributional assumptions.
+While their approach employs a bucketing-based sorting algorithm similar to ours, there are several key differences between their method and ours.
+First, their algorithm relies on detailed prior knowledge about the distribution; it explicitly requires a parameter $\varkappa_{2}$, which represents the “learning possibility” of the distribution (see Definition 3.2 in (zeighami2024theoretical) for details).
+Since $\varkappa_{2}$ depends on $\rho_{1}$ and $\rho_{2}$, their algorithm cannot be executed without knowing the values of $\rho_{1}$ and $\rho_{2}$ (or at least their lower and upper bounds).
+In contrast, our algorithm does not require these values, making it more widely applicable.
+Second, they do not provide experimental results.
+Since estimating $\rho_{1}$ and $\rho_{2}$ from real data is challenging, empirical evaluation of their method is difficult.
+On the other hand, since our algorithm does not depend on these specific values, it is easy to implement and evaluate experimentally.
+Finally, their algorithm lacks a worst-case complexity guarantee.
+In particular, there are cases where the algorithm may not terminate, making it impossible to give an upper bound on its worst-case complexity.
+In contrast, we provide a formal worst-case complexity guarantee.
+Limitations of the Current Theoretical Framework.
+Our proof of the $\mathcal{O}(n\log\log n)$ expected complexity (Theorem˜3.6), does not extend to distributions with $f(x)=0$ or $\infty$.
+A similar limitation is observed in Learned Indexes (zeighami2023distribution), and extending the theory to cover a broader class of distributions remains an open direction in both Learned Indexes and Learned Sorts contexts.
+One promising direction is to integrate more advanced CDF approximation methods with theoretical guarantees into our Learned Sort algorithm.
+By adopting a refined CDF model and a bucketing algorithm that satisfies the conditions of Lemma˜3.3, it may be possible to achieve stronger theoretical guarantees.
+Implementation Considerations and Optimizations.
+A straightforward implementation of PCF Learned Sort is not in-place, as it requires an auxiliary buffer for buckets nearly as large as the input.
+Engineering techniques from highly optimized sample-sort implementations (axtmann2017; axtmann2022engineering) suggest in-place variants without changing the asymptotic structure of our algorithm.
+Another important aspect is real-world efficiency: optimizing memory access patterns to enhance cache efficiency, reducing cache misses, and leveraging cache-aware data layouts could improve empirical performance without compromising theoretical guarantees.
+Dynamically tuning parameters such as $\alpha$, $\beta$, $\gamma$, and $\delta$ based on the input distribution may yield further performance gains while maintaining guarantees.
+Parallelization Potential.
+Although our analysis targets the sequential setting, the structure of PCF Learned Sort admits parallelization at several stages: training the CDF model, computing bucket IDs, scattering into buckets, and sorting buckets are all amenable to data parallelism.
+This observation is consistent with prior engineering work that integrates learned partitioning with high-performance (in-place) sample-sort pipelines and parallel learned-sorting frameworks (axtmann2017; carvalho2022towards; carvalho2023learnedsort).
+These results complement our theoretical analysis and suggest opportunities for developing parallel learned-sorting algorithms with provable guarantees.
 6 Conclusion
-We proposed PCF Learned Sort and proved theoretically that its worst-case computational complexity is $\mathcal{O}(n\log n)$ (without assumptions) and its expected computational complexity is $\mathcal{O}(n\log\log n)$ under mild assumptions on the distribution.
-We then confirm this computational complexity experimentally on both synthetic and real data.
-Furthermore, we provide a theoretical and intuitive explanation for the slow speed of Learned Sort on some datasets.
-This is the first study to theoretically support the experimental success of Learned Sort and provides insight into why and when Learned Sort is fast.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+We proposed PCF Learned Sort, which provides guarantees on both its expected and worst-case complexities.
+We then confirmed these computational complexities empirically on both synthetic and real data.
+This is the first study to support the empirical success of Learned Sort theoretically and provides insight into why Learned Sort is fast.
 Appendix A Proofs
-In this Appendix, we give the proofs omitted in the main paper.
-In Section A.1, we give the proof of Lemma 3.1, which is important for proving the worst-case complexity of PCF Learned Sort.
-Section A.2 and Section A.3 give proofs of Lemma 3.3 and Lemma 3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort.
-A.1 Proof of Lemma 3.1
+Here, we give the proofs omitted in the main paper.
+In Section˜A.1, we give the proof of Lemma˜3.1, which is important for proving the worst-case complexity of PCF Learned Sort.
+Section˜A.2 and Section˜A.3 give proofs of Lemma˜3.3 and Lemma˜3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort.
+The quantization-aware version of Lemma˜3.4 and Theorem˜3.6 is defined and the proof is given in the Section˜A.4.
+Finally, we provide the proofs of the theorems guaranteeing the complexity of Spline Learned Sort in Section˜A.5.
+A.1 Proof of Lemma˜3.1
 Proof.
-Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.1 and $\delta=\lfloor n^{d}\rfloor$.
-Let $S(n)$ be the worst-case complexity of the “standard” sort algorithm and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences).
-Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, and $\gamma+1=\mathcal{O}(n)$,
-$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n.$ (13)
+Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma˜3.1 and $\delta=\lfloor n^{d}\rfloor$.
+Let $S(n)$ be the worst-case complexity of the standard sort and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences).
+Since $S(n)=\mathcal{O}(nU(n))$ and the standard sort terminates after a finite number of operations,
+$\exists~C_{1},l_{1}~(>0),~~n\geq 0\;\;\Rightarrow\;\;S(n)\leq C_{1}+l_{1}nU(n).$ (3)
+Since $R(n)=\mathcal{O}(n)$ and $\gamma+1=\mathcal{O}(n)$,
 $\exists~n_{2},l_{2}~(>0),$ $~~n\geq n_{2}\;\;\Rightarrow\;\;R(n)\leq l_{2}n.$ (4)
 $\exists~n_{3},l_{3}~(>0),$ $~~n\geq n_{3}\;\;\Rightarrow\;\;\gamma+1\leq l_{3}n.$ (5)
-First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $P(n)\leq C$.
+In the following, we prove $P(n)=\mathcal{O}(nU(n)+n\log\log n)$ by mathematical induction.
+First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant C(>0)C~(>0) such that $P(n)\leq C$.
 That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
-This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the “standard” sort algorithm.
-Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$.
+This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the standard sort.
+Next, assume that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $P(n)\leq C+l(nU(n)+n\log\log n)$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$.
+Without loss of generality, we assume that $C\geq C_{1}$ and $l\geq l_{1}$.
 Let $\mathcal{S}_{\gamma,k}$ be the set consisting of all $(\gamma+1)$-dimensional vectors of positive integers whose sum is $k$, i.e., $\mathcal{S}_{\gamma,k}\coloneq\left\{\bm{s}\in\mathbb{Z}_{\geq 0}^{\gamma}\mid\sum_{i=1}^{\gamma+1}s_{i}=k\right\}$.
 Then,
-$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+ls_{i}\log s_{i}\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\left(C+ls_{i}\log k^{d}\right)\right\}\\ &\leq l_{2}k+C_{1}(\gamma+1)+l_{1}k\log k+C(\gamma+1)+lk\log k^{d}\\ &\leq l_{2}k+C_{1}l_{3}k+l_{1}k\log k+Cl_{3}k+lkd\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}-l(1-d)\right\}k\log k+(C+lk\log k).\end{split}$ (16)
+$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}U(s_{i})\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+l(s_{i}U(s_{i})+s_{i}\log\log s_{i})\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left(C+ls_{i}U(s_{i})+ls_{i}\log\log k^{d}\right)\\ &\leq l_{2}k+C(\gamma+1)+lkU(k)+lk\log\log k^{d}\\ &\leq l_{2}k+Cl_{3}k+lkU(k)+lk\log\log k+lk\log d\\ &\leq\left(l_{2}+Cl_{3}-l\log\frac{1}{d}\right)k+(C+l(kU(k)+k\log\log k)).\end{split}$ (6)
 Therefore, if we take $l$ such that
-$\frac{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}}{1-d}\leq l,$ (17)
-then $P(k)\leq C+lk\log k$ (note that the left side of the following equation is a constant independent of $k$).
-Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n\in\mathbb{N}$. ∎
-A.2 Proof of Lemma 3.3
+$\frac{l_{2}+Cl_{3}}{\log\frac{1}{d}}\leq l,$ (7)
+then $P(k)\leq C+l(kU(k)+k\log\log k)$ (note that the left side of Equation˜7 is a constant independent of $k$).
+Hence, by mathematical induction, it is proved that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $P(n)\leq C+l(kU(k)+lk\log\log k)$ for all $n\in\mathbb{N}$. ∎
+A.2 Proof of Lemma˜3.3
 Proof.
-The proof approach is the same as in Lemma 3.1, but in Lemma 3.3 the “expected” computational complexity is bounded.
-The following two randomnesses are considered to compute the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
-Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.3 and $\delta=\lfloor n^{d}\rfloor$.
-Let $S(n)$ be the expected complexity of the “standard” sort algorithm and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences).
-Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
-$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n,$ (18)
+The proof approach is the same as in Lemma˜3.1, but in Lemma˜3.3, the “expected” computational complexity is bounded.
+The following two randomnesses are considered for computing the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
+Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma˜3.3 and $\delta=\lfloor n^{d}\rfloor$.
+Let $S(n)$ be the expected complexity of the standard sort and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences).
+Since $S(n)=\mathcal{O}(n\log n)$ and the standard sort terminates after a finite number of operations,
+$\exists~C_{1},l_{1}~(>0),~~n\geq 0\;\;\Rightarrow\;\;S(n)\leq C_{1}+l_{1}n\log n.$ (8)
+Since $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
 $\exists~n_{2},l_{2}~(>0),$ $~~n\geq n_{2}\;\;\Rightarrow\;\;R(n)\leq l_{2}n,$ (9)
@@
 In the following, we prove $T(n)=\mathcal{O}(n\log\log n)$ by mathematical induction.
-First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $T(n)\leq C$.
+First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant C(>0)C~(>0) such that $T(n)\leq C$.
 That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
-Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
+Next, assume that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
 $\begin{split}T(k)&\leq R(k)+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot S(|\bm{c}_{j}|)+\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr\left[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor\right]\cdot\mathbb{E}\left[\sum_{j=1}^{\gamma+1}S(|\bm{c}_{j}|)\right]+\mathbb{E}\left[\sum_{j=1}^{\gamma+1}\mathbbm{1}[|\bm{c}_{j}|<\lfloor k^{d}\rfloor]\cdot T(|\bm{c}_{j}|)\right]\\ &\leq l_{2}k+\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor k^{d}\rfloor]\cdot\{C_{1}(\gamma+1)+l_{1}k\log k\}+\mathbb{E}\left[\sum_{i=1}^{\gamma+1}T(\min(\lfloor k^{d}\rfloor,|\bm{c}_{j}|))\right].\end{split}$ (12)
@@
 $\frac{l_{2}+C_{1}l_{3}l_{4}+l_{1}l_{3}+Cl_{4}}{\log\frac{1}{d}}\leq l,$ (16)
-then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation 26 is a constant independent of $k$).
-Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
-A.3 Proof of Lemma 3.4
-To prove Lemma 3.4, we first present the following lemma.
+then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation˜16 is a constant independent of $k$).
+Hence, by mathematical induction, it is proved that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
+A.3 Proof of Lemma˜3.4
+We first present the following lemma to prove Lemma˜3.4.
 Lemma A.1.
 Let $\bm{e}~(\in\mathcal{I}^{n})$ be a sorted version of $\bm{x}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ and $\Delta\coloneqq(x_{\mathrm{max}}-x_{\mathrm{min}})/\beta$ (where $x_{\mathrm{min}}$ and $x_{\mathrm{max}}$ are the minimum and maximum values of $\bm{x}_{\mathcal{I}}$, respectively).
-Also, define the set $\mathcal{S}_{r},\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
-$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (27)
+Also, define the set $\mathcal{S}_{r}$ and $\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
+$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~~~~~\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (17)
 Using this definition, define $Y_{jr},Z_{jr},Y_{r},Z_{r}$ as follows ($j=1,\dots,\alpha,~r=1,\dots,n$):
-$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~{}~{}~{}~{}~{}~{}~{}~{}Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},$ (28)
-$Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~{}~{}~{}~{}~{}~{}~{}~{}Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}.$ (29)
-If the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, then the following holds:
+$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~~~~~Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}$ (18)
+When using $\mathcal{M}_{\mathrm{PCF}}$, if the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, the following holds:
 $\left(r\geq\frac{\delta}{2}+1~\land~Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~\lor~\left(r\leq n-\frac{\delta}{2}~\land~Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right).$ (19)
@@
 ∎
-Using Lemma A.1, we can prove Lemma 3.4.
+Using Lemma˜A.1, we can prove Lemma˜3.4.
 Proof.
-[…]
-[…]
+Let $q=\max\limits_{y}\int_{y}^{y+\Delta}f_{\mathcal{I}}(x)dx$ (where $y$ is a value such that $(y,y+\Delta)\subseteq\mathcal{I}$).
+Then, from $\sigma_{1}\leq f(x)\leq\sigma_{2}$ for all $x\in\mathcal{I}$,
+$\begin{split}q&\leq\frac{\max_{y}\int_{y}^{y+\Delta}f(y)dy}{\int_{\mathcal{I}}f(x)dx}\\ &\leq\frac{\max_{y}\int_{y}^{y+\Delta}\sigma_{2}dy}{\int_{\mathcal{I}}\sigma_{1}dx}\\ &\leq\frac{\sigma_{2}\Delta}{\sigma_{1}(x_{\mathrm{max}}-x_{\mathrm{min}})}\\ &=\frac{\sigma_{2}}{\sigma_{1}\beta}.\end{split}$ (23)
 Thus, when $r\geq\frac{\delta}{2}+1$,
@@
 Here, when $K\geq 1$, we have
-$10\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<$ 1 . (37)
+$0\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<1.$ (26)
 Therefore, from the Chernoff bound,
@@
 $\Pr\left[Z_{r}\leq\frac{\alpha}{\gamma}\right]\leq\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (28)
-Thus, by defining $E_{r}$ to be the event “$e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma A.1,
+Thus, by defining $E_{r}$ to be the event $e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma˜A.1,
 $\begin{split}\Pr[E_{r}]&\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~\land~Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~\lor~\left(r\leq n-\frac{\delta}{2}~\land~Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\Pr\left[\left(r\geq\frac{\delta}{2}+1~\land~Y_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]+\Pr\left[\left(r\leq n-\frac{\delta}{2}~\land~Z_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)\right]\\ &\leq\begin{cases}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&\left(r<\frac{\delta}{2}+1~\lor~r>n-\frac{\delta}{2}\right)\\ 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},&(\mathrm{else})\end{cases}\\ &\leq 2\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (29)
@@
 $\begin{split}\Pr[\exists j,|\bm{c}_{j}|>\delta]&=\Pr[|\{j\mid|\bm{c}_{j}|>\delta\}|\geq 1]\\ &\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.\end{split}$ (32)
+A.4 The Quantization-Aware Version of Lemma˜3.4 and Theorem˜3.6
+In general, computers represent numbers in a finite number of bits, so the numbers they handle are inherently discrete.
+However, Theorem˜3.6, which states that the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$, does not cover discrete distributions.
+Here, we define the quantization process by which a computer represents numbers in finite bits and then show that the expected computational complexity of PCF Learned Sort is still $\mathcal{O}(n\log\log n)$, under the assumption that “the quantization is fine enough.”
+First, we assume the sampling and quantization process is as follows:
+Assumption A.2.
+For a range of values $\mathcal{D}~(\subseteq\mathbb{R})$, define $m~(\in\mathbb{N})$ contiguous regions $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{m}$ such that (i) they are disjoint from each other and (ii) together they form $\mathcal{D}$.
+For each region, determine representative values $r_{1},r_{2},\dots,r_{m}$.
+Here, $r_{1},r_{2},\dots,r_{m}$ are values contained in $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{m}$, respectively.
+For a value $x~(\in\mathcal{D})$, the quantized value of $x$, $x^{\prime}$, is obtained as $x^{\prime}=r_{i}$, where $i$ is the (only) $i$ such that $x\in\mathcal{D}_{i}$.
+The value $x$ is sampled according to the probability density function $f(x)$, but a computer keeps $x^{\prime}$ (instead of $x$) in $\log_{2}m$ bits with some quantization error.
+Also, for the interval $\mathcal{I}\subseteq\mathcal{D}$, we define $\eta(\mathcal{I})$ as follows.
+Definition A.3.
+$\eta(\mathcal{I})$ is the maximum width of $\mathcal{D}_{i}$ that intersects with interval $\mathcal{I}$, i.e.,
+$\eta(\mathcal{I})\coloneq\max\left\{|\mathcal{D}_{i}|\mid\mathcal{D}_{i}\cup\mathcal{I}\neq\varnothing,i=1,\dots,m\right\},$ (33)
+where $|\mathcal{D}_{i}|$ is the width of the range $\mathcal{D}_{i}$.
+We can prove that $\eta(\mathcal{I})$ is the upper bound of the quantization error of $x$ in $\mathcal{I}$, i.e., $|x-x^{\prime}|\leq\eta(\mathcal{I})$ when $x\in\mathcal{I}$.
+Now, the assumption that the quantization is “fine enough” is specifically defined as follows.
+Assumption A.4.
+(Recall that our PCF Learned Sort recursively calls its own algorithm.
+Each time of recursion, the range of values of interest $\mathcal{I}$ changes, and the length of the array of interest $n$ also changes.) For all $\mathcal{I}$ and $n$ that appear in the algorithm, the following holds:
+$\frac{\beta\eta(\mathcal{I})}{|\mathcal{I}|}\leq\frac{1}{2}.$ (34)
+We show intuitively and empirically that this is a satisfactory assumption.
+First, to show intuitively that this assumption is satisfactory, we give an example.
+Let $\mathcal{I}$ be a range of values that can be represented by a 64-bit double (1 bit for the sign, 11 bits for the exponent part, and 52 bits for the mantissa part), that is, $\mathcal{I}=[-1.79\times 10^{308},1.79\times 10^{308}]$.
+The quantization is performed by mapping each value to the nearest number that can be represented by a 64-bit double.
+In this setting, $\eta(\mathcal{I})=1.99\times 10^{292}$, $|\mathcal{I}|=3.59\times 10^{308}$.
+$\eta(\mathcal{I})/|\mathcal{I}|=5.55\times 10^{-17}$.
+Thus, for usual $\beta$, $\beta\leq 9\times 10^{15}$, Equation˜34 holds.
+Second, we show that Equation˜34 is a satisfactory assumption empirically.
+In the 1,280 measurements, where 10 measurements each for 16 different $n(\in\{10^{3},2\times 10^{3},5\times 10^{3},\dots,10^{8}\})$ on 8 different datasets, $5.23\times 10^{7}$ pairs of $(\mathcal{I},\beta)$ appeared (we set $\beta=\left\lfloor n^{3/4}\right\rfloor$), and the left side of Equation˜34 is at most $8.11\times 10^{-9}$, indicating that Equation˜34 is always true with a margin.
+Under this definition and assumption about quantization, we can prove the quantization-aware version of Lemma˜3.4.
+Lemma A.5 (Quantization-aware version of Lemma˜3.4).
+Let $\sigma_{1}$ and $\sigma_{2}$ be respectively the lower and upper bounds of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
+That is, $x\in\mathcal{D}\;\;\Rightarrow\;\;\sigma_{1}\leq f(x)\leq\sigma_{2}$.
+Also, let $\bm{x}^{\prime}_{\mathcal{I}}$ be the array created by the quantization of $\bm{x}~(\in\mathcal{I}^{n})$ in the manner defined in ˜A.2.
+Then, in model-based bucketing of $\bm{x}^{\prime}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~(\subseteq\mathcal{D})$ under ˜A.4:
+$K\coloneqq\frac{\gamma\delta}{2n}-\frac{4\sigma_{2}\gamma}{\sigma_{1}\beta}\geq 1\;\;\Rightarrow\;\;\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K^{\prime}}{2\gamma}\left(1-\frac{1}{K^{\prime}}\right)^{2}\right\}.$ (35)
+The proof of Lemma˜A.5 is done in the same way as Lemma˜3.4. That is, we first prove the following lemma.
+Lemma A.6 (Quantization-aware version of Lemma˜A.6).
+Let $\bm{e}^{\prime}~(\in\mathcal{I}^{n})$ be a sorted version of $\bm{x}^{\prime}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ and $\Delta\coloneqq(x^{\prime}_{\mathrm{max}}-x^{\prime}_{\mathrm{min}})/\beta$ (where $x^{\prime}_{\mathrm{min}}$ and $x^{\prime}_{\mathrm{max}}$ are the minimum and maximum values of $\bm{x}^{\prime}_{\mathcal{I}}$, respectively).
+Also, define the set $\mathcal{S}^{\prime}_{r},\mathcal{T}^{\prime}_{r}$ as follows ($r=1,\dots,n$):
+$\mathcal{S}^{\prime}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta+2\eta(\mathcal{I})<e_{k}\leq e_{r}-\Delta-2\eta(\mathcal{I})\},$ (36)
+$\mathcal{T}^{\prime}_{r}=\{k\mid e_{r}+\Delta+2\eta(\mathcal{I})\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta-2\eta(\mathcal{I})\}.$ (37)
+Using this definition, define $Y^{\prime}_{jr},Z^{\prime}_{jr},Y^{\prime}_{r},Z^{\prime}_{r}$ as follows ($j=1,\dots,\alpha,~r=1,\dots,n$):
+$Y^{\prime}_{jr}=\begin{cases}1&(j\in\mathcal{S}^{\prime}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Z^{\prime}_{jr}=\begin{cases}1&(j\in\mathcal{T}^{\prime}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Y^{\prime}_{r}=\sum_{j=1}^{\alpha}Y^{\prime}_{jr},~~~~~Z^{\prime}_{r}=\sum_{j=1}^{\alpha}Z^{\prime}_{jr}.$ (38)
+If the size of the bucket to which $e^{\prime}_{r}$ is allocated is greater than or equal to $\delta$, then the following holds:
+$\left(r\geq\frac{\delta}{2}+1~\land~Y^{\prime}_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~\lor~\left(r\leq n-\frac{\delta}{2}~\land~Z^{\prime}_{r}\leq\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right).$ (39)
+Proof.
+The proof method is exactly the same as for Lemma˜A.1. That is, by taking the difference between the IDs of the buckets to which $e^{\prime}_{r+\delta/2}$ and $e^{\prime}_{r}$ are assigned,
+$\begin{split}&\left(\lfloor\tilde{F}(e^{\prime}_{r+\delta/2})\gamma\rfloor+1\right)-\left(\lfloor\tilde{F}(e^{\prime}_{r})\gamma\rfloor+1\right)\\ >&\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta-2\eta(\mathcal{I})\}\right|-\left(\frac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta+2\eta(\mathcal{I})\}\right|+1\right)\\ =&\frac{\gamma}{\alpha}\left|\{j\mid e_{r}+\Delta+2\eta(\mathcal{I})<a_{j}\leq e_{r+\delta/2}-\Delta-2\eta(\mathcal{I})\}\right|-1\\ =&\frac{\gamma}{\alpha}Z^{\prime}_{r}-1\\ \geq&0,\end{split}$ (40)
+when $Z^{\prime}_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor$. Thus, we can prove that when
+$\left(r<\frac{\delta}{2}+1~\lor~Y^{\prime}_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)~\land~\left(r>n-\frac{\delta}{2}~\lor~Z^{\prime}_{r}>\left\lfloor\frac{\alpha}{\gamma}\right\rfloor\right)$ (41)
+holds, $e^{\prime}_{r+\delta/2}$ and $e^{\prime}_{r}$ are assigned to the different bucket and $e^{\prime}_{r-\delta/2}$ and $e^{\prime}_{r}$ are assigned to the different bucket.
+Then, the contraposition of the lemma is proved.
+∎
+Using Lemma˜A.6, we can prove Lemma˜A.5.
+Proof.
+Let $q^{\prime}=\max\limits_{y}\int_{y}^{y+\Delta 2\eta(\mathcal{I})}f_{\mathcal{I}}(x)dx$ (where $y$ is a value such that $(y,y+\Delta+2\eta(\mathcal{I})\subseteq\mathcal{I}$).
+Then, from $\sigma_{1}\leq f(x)\leq\sigma_{2}$ for all $x\in\mathcal{I}$,
+$\begin{split}q^{\prime}&\leq\frac{\max_{y}\int_{y}^{y+\Delta+2\eta(\mathcal{I})}f(y)dy}{\int_{\mathcal{I}}f(x)dx}\\ &\leq\frac{\max_{y}\int_{y}^{y+\Delta+2\eta(\mathcal{I})}\sigma_{2}dy}{\int_{\mathcal{I}}\sigma_{1}dx}\\ &=\frac{\sigma_{2}(\Delta+2\eta(\mathcal{I}))}{\sigma_{1}(x^{\prime}_{\mathrm{max}}-x^{\prime}_{\mathrm{min}})}\\ &\leq\frac{\sigma_{2}}{\sigma_{1}\beta}\cdot\left(1+\frac{2\beta\eta(\mathcal{I})}{\left|\mathcal{I}\right|}\right)\\ &\leq\frac{2\sigma_{2}}{\sigma_{1}\beta}.\end{split}$ (42)
+The last inequality is obtained by ˜A.4.
+Thus, when $r\geq\frac{\delta}{2}+1$,
+$\begin{split}\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}^{\prime}_{r}\right|\right]&=\mathbb{E}\left[\frac{\delta}{2}-\left|\{k\mid e_{r-\delta/2}+\Delta+2\eta(\mathcal{I})<e_{k}\leq e_{r}-\Delta-2\eta(\mathcal{I})\}\right|\right]\\ &=\mathbb{E}\left[\left|\{k\mid e_{r-\delta/2}<e_{k}\leq e_{r-\delta/2}+\Delta+2\eta(\mathcal{I})\}\right|\right]+\mathbb{E}\left[\left|\{k\mid e_{r}-\Delta-2\eta(\mathcal{I})<e_{k}\leq e_{r}\}\right|\right]\\ &\leq nq^{\prime}+nq^{\prime}\\ &\leq\frac{4\sigma_{2}n}{\sigma_{1}\beta}.\end{split}$ (43)
+Thus, when $r\geq\frac{\delta}{2}+1$,
+$\begin{split}\mathbb{E}[Y^{\prime}_{r}]&=\frac{\alpha}{n}\mathbb{E}\left[\left|\mathcal{S}^{\prime}_{r}\right|\right]\\ &=\frac{\alpha}{n}\left(\frac{\delta}{2}-\mathbb{E}\left[\frac{\delta}{2}-\left|\mathcal{S}^{\prime}_{r}\right|\right]\right)\\ &\geq\frac{\alpha\delta}{2n}-\frac{4\sigma_{2}\alpha}{\sigma_{1}\beta}\\ &=\frac{\alpha K^{\prime}}{\gamma}.\end{split}$ (44)
+From this point forward, by proceeding in exactly the same way as the proof of Lemma˜3.4, we can prove the following using Lemma˜A.1:
+$K^{\prime}\geq 1\;\;\Rightarrow\;\;\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K^{\prime}}{2\gamma}\left(1-\frac{1}{K^{\prime}}\right)^{2}\right\}.$ (45)
+Using Lemma˜A.5, we can prove the following theorem.
+Theorem A.7 (Quantization-aware version of Theorem˜3.6).
+Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
+Also, assume that the input array is quantized in a way that satisfies ˜A.2 and ˜A.4.
+Then, the expected complexity of PCF Learned Sort with $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ is $\mathcal{O}(n\log\log n)$.
+Proof.
+When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$.
+Since $K^{\prime}=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K^{\prime}\geq 1$ for sufficiently large $n$, and
+$\frac{2n}{\delta}\exp\left\{-\frac{\alpha K^{\prime}}{2\gamma}\left(1-\frac{1}{K^{\prime}}\right)^{2}\right\}=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\leq\mathcal{O}\left(\frac{1}{\log n}\right).$ (46)
+Therefore, from Lemma˜3.3 and Lemma˜A.5, the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
+A.5 Proofs for Spline Learned Sort
+Definition of Spline Learned Sort.
+We now formally define Spline Learned Sort, which we propose as a variant of PCF Learned Sort.
+In particular, Spline Learned Sort replaces $\mathcal{M}_{\mathrm{PCF}}$—the bucketing method that uses PCF as the CDF model—with $\mathcal{M}_{\mathrm{Spline}}$, a bucketing method based on a spline-based CDF model.
+Below, we give a rigorous description of $\mathcal{M}_{\mathrm{Spline}}$.
+The training algorithm of the CDF model in $\mathcal{M}_{\mathrm{Spline}}$ is the same as in $\mathcal{M}_{\mathrm{PCF}}$.
+That is, the parameters $\alpha,\beta\in\mathbb{N}$ are determined by $n$.
+Then, following this setting, we sample an array $\bm{a}$ of length $\alpha$ from the input array $\bm{x}$, which is subsequently bin-counted (as described in Section˜3.3).
+For later explanation and proofs, we introduce additional terminology.
+We define the vector of thresholds used for counting as $\bm{t}\in\mathbb{R}^{\beta+2}$: $t_{i}=x_{\mathrm{min}}+(x_{\mathrm{max}}-x_{\mathrm{min}})(i-1)/\beta$ for $i=1,2,\dots,\beta+2$.
+Thus, the $i$-th bin during CDF model training corresponds to the interval $t_{i},t_{i+1})$.
+Accordingly, the non-decreasing, non-negative array $\bm{b}$ formed by bin-counting is given by $b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid a_{j}<t_{i+1}\right\}\right|$ for $i=1,2,\dots,\beta+1$.
+We also denote by $F_{\bm{a}}(x)$ the empirical CDF of the sampled array $\bm{a}$: $F_{\bm{a}}(x)=\left|\{j\in\{1,2,\dots,\alpha\}\mid a_{j}\leq x\}\right|/\alpha$.
+The inference algorithm of the CDF model in $\mathcal{M}_{\mathrm{Spline}}$ is similar to but slightly different from that in $\mathcal{M}_{\mathrm{PCF}}$.
+In PCF, the prediction is constant within each interval $t_{i},t_{i+1})$.
+In contrast, in the spline-based case, the prediction is obtained by linearly interpolating the empirical CDF values at the endpoints of the interval.
+Specifically, the output $\tilde{F}(x)$ for input $x$ is obtained as follows.
+As in PCF, we compute $i(x)=\left\lfloor\frac{(x-x_{\min})\beta}{x_{\max}-x_{\min}}\right\rfloor+1$, so that $x\in[t_{i(x)},t_{i(x)+1})$.
+While PCF returns $F_{\bm{a}}(t_{i(x)+1})=b_{i(x)}/\alpha$ as the CDF prediction, the spline-based method linearly interpolates between $F_{\bm{a}}(t_{i(x)})$ and $F_{\bm{a}}(t_{i(x)+1})$:
+$\tilde{F}(x)$ $=F_{\bm{a}}(t_{i(x)})+\frac{x-t_{i(x)}}{t_{i(x)+1}-t_{i(x)}}\cdot(F_{\bm{a}}(t_{i(x)+1})-F_{\bm{a}}(t_{i(x)}))$ (47)
+$=\frac{1}{\alpha}\left(b_{i(x)-1}+\frac{x-t_{i(x)}}{t_{i(x)+1}-t_{i(x)}}\cdot(b_{i(x)}-b_{i(x)-1})\right),$ (48)
+with $b_{0}=0$.
+Unlike PCF, which only considers which bin a value falls into, the spline-based method also accounts for the relative position of the value within the bin to compute the predicted CDF $\tilde{F}(x)$.
+The asymptotic computational complexity for training and inference is the same as PCF: $\mathcal{O}(n)$ for training and $\mathcal{O}(1)$ per element for inference.
+However, unlike PCF (which only requires array lookups), the spline-based approach requires additional subtractions and multiplications during inference, leading to a larger constant factor in runtime.
+Theorems Guaranteeing the Complexity of Spline Learned Sort.
+When $\mathcal{M}_{\mathrm{Spline}}$ is used as the bucketing method in our learned sort framework, the same complexity guarantees as in PCF Learned Sort hold.
+In particular, we obtain the following worst-case guarantee (Theorem A.8) and expected-case guarantee (Theorem A.9).
+Theorem A.8 (Spline version of Theorem˜3.5).
+If $\mathcal{M}_{\mathrm{Spline}}$ is the bucketing method, the worst-case complexity of the standard sort is $\mathcal{O}(nU(n))$ (where $U(n)$ is a non-decreasing function), and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the worst-case complexity of Spline Learned Sort is $\mathcal{O}(nU(n)+n\log\log n)$.
+Proof.
+The proof is identical to that of Theorem˜3.5, using Lemma˜3.1. ∎
+Theorem A.9 (Spline version of Theorem˜3.6).
+Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density $f(x)$ on $\mathcal{D}$, and assume $0<\sigma_{1}\leq\sigma_{2}<\infty$.
+If $\mathcal{M}_{\mathrm{Spline}}$ is the bucketing method, the expected complexity of the standard sort is $\mathcal{O}(n\log n)$, and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the expected complexity of Spline Learned Sort is $\mathcal{O}(n\log\log n)$.
+To prove this theorem, we first establish the following lemma.
+Lemma A.10 (Spline version of Lemma˜A.1).
+Define $\bm{e},\mathcal{S}_{r},\mathcal{T}_{r},Y_{jr},Z_{jr},Y_{r},Z_{r}$ exactly as in Lemma˜A.1 ($j=1,\dots,\alpha,~r=1,\dots,n$).
+That is, let $\bm{e}\in\mathcal{I}^{n}$ be the sorted version of $\bm{x}_{\mathcal{I}}\in\mathcal{I}^{n}$ and set $\Delta\coloneqq(x_{\mathrm{max}}-x_{\mathrm{min}})/\beta$ (where $x_{\mathrm{min}}$ and $x_{\mathrm{max}}$ are the minimum and maximum values of $\bm{x}_{\mathcal{I}}$, respectively).
+Also define
+$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~~~~~\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\},$ (49)
+and
+$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{otherwise})\end{cases},~~~~~Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{otherwise})\end{cases},~~~~~Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~~~~~Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}.$ (50)
+When using $\mathcal{M}_{\mathrm{Spline}}$, if the size of the bucket to which $e_{r}$ is assigned is at least $\delta$, then the same condition as in Equation˜19 holds; namely,
+$\left(r\geq\tfrac{\delta}{2}+1~\land~Y_{r}\leq\left\lfloor\tfrac{\alpha}{\gamma}\right\rfloor\right)~\lor~\left(r\leq n-\tfrac{\delta}{2}~\land~Z_{r}\leq\left\lfloor\tfrac{\alpha}{\gamma}\right\rfloor\right).$ (51)
+Proof.
+The proof proceeds exactly as in Lemma˜A.1, by contraposition.
+That is, assume $\left(r<\tfrac{\delta}{2}+1~\lor~Y_{r}>\lfloor\tfrac{\alpha}{\gamma}\rfloor\right)~\land~\left(r>n-\tfrac{\delta}{2}~\lor~Z_{r}>\lfloor\tfrac{\alpha}{\gamma}\rfloor\right)$, and prove that $e_{r}$ is assigned to a bucket smaller than $\delta$.
+First, we show that $e_{r}$ and $e_{\min(r+\delta/2,n+1)}$ are assigned to different buckets.
+As in the proof of Lemma˜A.1, the case $n-\delta/2<r\leq n$ is immediate; hence we focus on $r\leq n-\delta/2$.
+The bucket ID of $e_{r}$ is $\lfloor\tilde{F}(e_{r})\gamma\rfloor+1$, while that of $e_{r+\delta/2}$ is $\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1$.
+By the definition of the spline-based CDF model and the inequalities $t_{i(e_{r})+1}\leq e_{r}+\Delta,~t_{i(e_{r+\delta/2})}\geq e_{r+\delta/2}-\Delta$, we obtain
+$\tilde{F}(e_{r})$ $\leq F_{\bm{a}}(t_{i(e_{r})+1})=\tfrac{1}{\alpha}\left|\{j\mid a_{j}\leq t_{i(e_{r})+1}\}\right|\leq\tfrac{1}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|,$ (52)
+$\tilde{F}(e_{r+\delta/2})$ $\geq F_{\bm{a}}(t_{i(e_{r+\delta/2})})=\tfrac{1}{\alpha}\left|\{j\mid a_{j}\leq t_{i(e_{r+\delta/2})}\}\right|\geq\tfrac{1}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|.$ (53)
+Hence, their bucket IDs satisfy
+$\lfloor\tilde{F}(e_{r})\gamma\rfloor+1$ $\leq\tfrac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r}+\Delta\}\right|+1,$ (54)
+$\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1$ $>\tfrac{\gamma}{\alpha}\left|\{j\mid a_{j}\leq e_{r+\delta/2}-\Delta\}\right|,$ (55)
+which are identical to Equations˜20 and 21. Thus,
+$\left(\lfloor\tilde{F}(e_{r+\delta/2})\gamma\rfloor+1\right)-\left(\lfloor\tilde{F}(e_{r})\gamma\rfloor+1\right)>0,$ (56)
+and hence $e_{r}$ and $e_{r+\delta/2}$ are assigned to different buckets.
+By a symmetric argument, $e_{\max(0,r-\delta/2)}$ and $e_{r}$ are also assigned to different buckets.
+Therefore, the bucket containing $e_{r}$ has size at most $\delta-1$, completing the contraposition.
+∎
+From Lemma A.10, we obtain the following theorem.
+Theorem A.11 (Spline version of Lemma˜3.4).
+Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds of the probability density $f(x)$ on $\mathcal{D}$, and assume $0<\sigma_{1}\leq\sigma_{2}<\infty$ (i.e., $\sigma_{1}\leq f(x)\leq\sigma_{2}$ for all $x\in\mathcal{D}$).
+Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}\in\mathcal{I}^{n}$ into $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{Spline}}$, the following holds for any interval $\mathcal{I}\subseteq\mathcal{D}$:
+$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}\geq 1\;\;\Rightarrow\;\;\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\tfrac{1}{K}\right)^{2}\right\}.$ (57)
+Proof.
+The proof is identical to that of Lemma˜3.4, using Lemma A.10. ∎
+Finally, using Theorem A.11, we can prove Theorem A.9.
+Proof of Theorem˜A.9.
+The proof follows exactly the same steps as Theorem˜3.6, applying Theorem A.11 to Lemma˜3.3. ∎
+Appendix B Real Dataset Details
+Table 1: Statistics of real-world datasets used in our experiments.
+Dataset Size # Unique Values % of Duplicates
+Chicago [Start] (chic) 39,588,772 216,610 99.45
+Chicago [Tot] (chic) 39,199,154 6,887 99.98
+NYC [Pickup] (nyc) 100,000,000 26,666,741 73.33
+NYC [Dist] (nyc) 200,107,656 2,060 100.00
+NYC [Tot] (nyc) 199,964,459 7,428 100.00
+SOF [Humidity] (sof) 96,318,228 8,128 99.99
+SOF [Pressure] (sof) 96,317,180 862,502 99.10
+SOF [Temperature] (sof) 96,432,856 5,638 99.99
+Wiki (sosd-vldb) 200,000,000 90,437,011 54.78
+OSM (sosd-vldb) 800,000,000 799,469,195 0.07
+Books (sosd-vldb) 800,000,000 799,994,961 0.00
+Face (sosd-vldb) 199,998,000 199,998,000 0.00
+Stocks [Volume] (stks) 27,596,686 325,654 98.82
+Stocks [Open] (stks) 27,596,686 830,758 96.99
+Stocks [Date] (stks) 27,596,686 14,646 99.95
+Stocks [Low] (stks) 27,596,686 863,701 96.87
+In the main text, we provided only a concise list of the real datasets used in our experiments. Here we describe each dataset in more detail.
+  • Chicago [Start, Tot]: The Chicago Taxi Trips dataset includes taxi trips reported to the City of Chicago in its role as a regulatory agency over the last six years.
+The data to be sorted includes trip starting timestamps and total fare amounts.
+  • NYC [Pickup, Dist, Tot]: The New York City yellow taxi trip dataset includes trip pickup datetimes, trip distances, and total fare amounts.
+  • SOF [Humidity, Pressure, Temperature]: The Sofia dataset contains time-series air quality metrics (humidity, pressure, and temperature) measured at 1-minute intervals from outdoor sensors in Sofia, Bulgaria.
+  • Wiki: The Wikipedia dataset contains article edit timestamps (sosd-vldb).
+  • OSM: Uniformly sampled OpenStreetMap locations represented as Google S2 CellIds (sosd-vldb).
+  • Books: Book sale popularity data from Amazon (sosd-vldb).
+  • Face: The FB dataset contains an upsampled set of Facebook user IDs obtained via random walks on the FB social graph (sosd-vldb).
+As in (kristo2021defeating; ferragina2025balancedsort), the outliers greater than the 0.99999 quantile are discarded.
+  • Stocks [Volume, Open, Date, Low]: The Stocks dataset contains historical daily opening, low prices, trading volumes, and dates for all NASDAQ tickers (stocks and ETFs), retrieved via the yfinance Python package up to April 1, 2020.
+Furthermore, the detailed statistics of these datasets, including their sizes, the number of unique values, and the percentage of duplicates, are summarized in Table˜1.
+The percentage of duplicates is computed as $100\times\left(1-\frac{|\,\text{unique\_values}\,|}{n}\right)$, where $n$ is the total number of elements and unique_values is the set of distinct elements.
+In particular, for several datasets (Chicago [Start, Tot], NYC [Dist, Tot], SOF [Humidity, Pressure, Temperature], and Stocks [Volume, Open, Date, Low]), the number of unique values is extremely small, accounting for less than 3.2% of the total elements.
+Appendix C Experiments in Adversarial Environments
+Figure 5: Time to sort the array in adversarial environments.
+Below each graph is a histogram that visualizes the distribution of each dataset.
+The standard deviation of the $10$ measurements is represented by the shaded area.
+To evaluate the robustness of our Learned Sort, we conducted experiments under adversarially constructed inputs that explicitly violate the assumptions of Lemma˜3.3.
+Specifically, given an original dataset of size $n$, we injected $n$ duplicate elements of a randomly chosen value inside the range $x_{\mathrm{min}},x_{\mathrm{max}}]$, where $x_{\mathrm{min}}$ and $x_{\mathrm{max}}$ is the minimun and maximum value of the input array $\bm{x}$.
+As a result, the array length doubled to $2n$, and the constructed distribution contained a point mass of probability one-half, leading to a huge probability density at that value.
+This setting intentionally breaks the assumption $\sigma_{2}<\infty$ required in Lemma˜3.3.
+Figure˜5 reports the sorting time on such adversarial datasets.
+The results show that the expected bound of $\mathcal{O}(n\log\log n)$ for our PCF Learend Sort no longer holds once the assumptions are violated, and the performance of PCF Learned Sort occasionally degrades to the level of std::sort.
+However, we consistently observed that PCF Learned Sort never exceeded the worst-case complexity of $\mathcal{O}(n\log n)$, and thus its runtime remained comparable to std::sort.
+On the other hand, we found that Learned Sort 2.0 (kristo2021defeating-github), which has $\mathcal{O}(n^{2})$ worst-case complexity, sometimes exhibited severe slowdowns.
+For example, on data of size $n=10^{5}$ sampled from a Normal distribution with injected duplicates, the sorting time of Learned Sort 2.0 reached up to 10.89 seconds, whereas PCF Learned Sort required at most 0.021 seconds.
+We also found that BLS, ULS, and Learned Sort 2.1 (ferragina2025balancedsort), which also have $\mathcal{O}(n^{2})$ worst-case complexity, sometimes exhibited substantial performance degradation.
+For example, on data of size $n=10^{6}$ sampled from a Normal distribution with injected point masses, Learned Sort 2.1 took up to 0.154 seconds, compared with 0.042 seconds for PCF Learned Sort and 0.078 seconds for std::sort.
+These results highlight the fragility of sorting algorithms without strong worst-case complexity guarantees, a limitation shared by many existing learned sorting algorithms.
+They underscore the importance of developing learned sorting algorithms—such as our PCF Learned Sort—that combine practical efficiency with rigorous worst-case complexity guarantees.
+Appendix D Runtime Profiling
+Figure 6: Breakdown of runtime components in PCF Learned Sort.
+The total runtime is decomposed into (i) model training, (ii) bucketing, (iii) standard sort applied to buckets smaller than $\tau$, and (iv) standard sort applied to buckets larger than $\delta$.
+Figure 7: Breakdown of runtime components in PCF Learned Sort in adversarial environments.
+The total runtime is decomposed into (i) model training, (ii) bucketing, (iii) standard sort applied to buckets smaller than $\tau$, and (iv) standard sort applied to buckets larger than $\delta$.
+To better understand the practical behavior of PCF Learned Sort, we profiled the runtime of its major components: (i) model training, (ii) bucketing, (iii) standard sort applied to buckets that are too small ($\tau$), and (iv) standard sort applied to buckets that are too large ($\geq\delta$).
+The results are shown in Figure˜6.
+For small input sizes $n$, the majority of the runtime is typically spent on applying the standard sort to buckets smaller than $\tau$.
+As $n$ increases, the relative cost shifts, and most of the runtime is instead consumed by the bucketing stage.
+Model training remains consistently lightweight, comparable to or faster than bucketing, and never exceeds roughly one quarter of the total runtime.
+Thus, training does not become the dominant factor in practice.
+Figure˜7 shows the results for adversarially clustered inputs (as described in Appendix˜C).
+Under the adversarial setting, the bucketing algorithm more frequently produces large buckets, which are then handled by standard sort.
+In these cases, a larger fraction of the runtime is attributed to this fallback mechanism.
+Overall, these analyses clarify which components of PCF Learned Sort dominate the runtime under different conditions.
+They also highlight that the algorithm adapts gracefully: training and bucketing scale well, while the worst-case fallback ensures robustness without catastrophic slowdowns.